### v2

# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

Abstract
We study the first passage times of discrete-time branching random walks in ${\mathbb{R}}^{d}$ where $d\geqslant 1$. Here, the genealogy of the particles follows a supercritical Galton-Watson process. We provide asymptotics of the first passage times to a ball of radius one at a distance $x$ from the origin, conditioned upon survival. We explicitly provide the linear dominating term and the logarithmic correction term as functions of $x$. The asymptotics are precise up to an order of $o_{\mathbb{P}}(\log x)$ for general jump distributions and up to $O_{\mathbb{P}}(\log\log x)$ for spherically symmetric jumps. A crucial ingredient of both results is the tightness of first passage times. We also discuss an extension of the first passage time analysis to a modified branching random walk model that has been proven to successfully capture shortest path statistics in polymer networks.
first passage time,
keywords:
[class=MSC]
keywords:
[…]
, , and
1 Introduction
We consider a discrete-time branching random walk (BRW) on $\mathbb{R}^{d}$, where particles in $\mathbb{R}^{d}$ perform independent random walks until they randomly branch into more particles or die. Spatial branching processes, including BRW and its continuous-time sibling branching Brownian motion (BBM), have a long history and a wide range of applications in ecology, population biology, and modeling epidemics ([ermakova2019branching, fisher1937wave, kolmogorov1937etude, konig2020branching, kot2004stochasticity, wang1980convergence]). The study of the extremal behavior of spatial branching processes has gained increasing attention over recent years. The mathematical techniques applied in those studies are related to developments in numerous other fields, e.g., discrete Gaussian free field [bramson2016convergence2, ding2014extreme], random unitary matrices [arguin2017maximum], Riemann zeta function [arguin2017maxima], random multiplicative functions [harper2020moments], spin glasses [bovier2017gaussian], holomorphic multiplicative chaos [gu2024universality], among many others that exhibit a log-correlated structure. We refer to [shi2015branching, zeitouni2016branching] for lecture notes on BRW and related topics.
The study of the extrema for one-dimensional BBM traces back to [bramson1978maximal], which established a precise asymptotic for the maximum as a function of time. For general branching random walks, [addario2009minima] provided a precise asymptotic in terms of the jump distribution of the process. Finer behavior near the frontier is also well understood since the works of [aidekon2013convergence, bramson2016convergence, lalley1987conditional, madaule2017convergence]. Nevertheless, less is known in dimensions greater than one. To name a few examples, there are studies of the maximum norm for BBM ([berestycki2024extremal, kim2023maximum, kim2024shape, mallein2015maximal, stasinski2021derivative]) and very recently of BRW ([bezborodov2023maximal]), as well as the range of spatial branching processes ([kyprianou2002asymptotic, zhang2024large]). These works in higher dimensions typically assumed that the increment distribution is spherically symmetric. A notable exception is [uchiyama1982spatial], which provided estimates on the number of particles in a linearly shifted ball in time for a continuous-time version of the BRW.
Our primary focus is on the first passage times (FPT) for branching random walks in $\mathbb{R}^{d}$. By definition, the FPT is the first time for particles to reach a prescribed region. For BBM, [zhang2024modeling] established a precise asymptotic formula (up to $O_{\mathbb{P}}(1)$) for the FPT; see (62) below. In this paper, we provide FPT asymptotics for BRW that may not have a spherically symmetric jump distribution (while the case of spherically symmetric jumps is of special interest for our applications; see Section 1.4). The main results are explained in Section 1.2 and their proofs are presented in Section 2. Applications to polymer physics will be briefly discussed in Section 1.4 and further detailed in Section 3.1.
We take slightly different routes depending on whether the jump distribution is spherically symmetric or not. In the symmetric case, a key part is establishing asymptotic upper and lower bounds on the number of particles that are near the frontier for a one-dimensional BRW (Proposition 8), where our approach follows the modified second moment method of [bramson2016convergence]. In the non-spherically symmetric case, our argument relies heavily on the theory of random walks in cones [denisov2015random]. In both settings, we prove the tightness around the median of the FPT (Theorem 2) using independent techniques, building upon ideas from [mcdiarmid1995minimal]. A more detailed proof sketch can be found in Section 1.3.
1.1 Notation and assumptions
Let us start with a formal definition of the branching random walk (BRW). We consider a discrete-time BRW model with offspring distribution $\{p_{i}\}_{i\geqslant 0}$. This means that at each time step, each particle is independently replaced by $i$ particles at the same location with probability $p_{i}$. The particles in each generation have i.i.d. displacements. The case of $i=0$ corresponds to the particle being terminated. Let $\rho=\sum_{i}i\,p_{i}$ be the mean of the offspring distribution. We will always assume the supercritical case $\rho>1$. The $d$-dimensional random walk increment is given by $\bm{\xi}$}, and the first particle starts from the origin $\mathbf{0}\in\mathbb{R}^{d}$. Typically, in this paper, a bold symbol refers to a vector.
Our goal is to study the asymptotics of the FPTs of the BRW. For $x\in\mathbb{R}$, we let $B_{x}$ denote a ball of radius one [^1] Replacing the radius one by any fixed positive constant does not change the analysis; the same proof goes through. centered at $(x,0,\dots,0)$ in $\mathbb{R}^{d}$. We let $V_{n}$ denote the collection (i.e., set) of particle locations at time step $n$, and $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ denotes the $d$-dimensional random walk that leads to $v\in V_{n}$. We define the FPT $\tau_{x}$ of the BRW to $B_{x}$ as
$\tau_{x}=\min\{n\geqslant 0:\exists\,v\in V_{n},\penalty 10000\ \bm{\eta}_{v,n}(n)\in B_{x}\}.$
In what follows, we will separately analyze the cases depending on whether the law of $\bm{\xi}$} is spherically symmetric or not. [^2] A probability distribution on $\mathbb{R}^{d}$ is spherically symmetric if it is invariant under any orthonormal transformation. Equivalent definitions can be found in Theorem 2.5 of [fang2018symmetric]. In both cases, we impose the following assumptions on the BRW:
  • (A1)
    the offspring distribution has a finite second moment, i.e., $\sum_{i}i^{2}\,p_{i}<\infty$;
  • (A2)
    the law of $\bm{\xi}$} is integrable and centered, i.e., $\mathbb{E}[{\bm{\xi}}]=\mathbf{0}$.
For a stochastic process $\{T_{x}\}_{x\geqslant 0}$ and a positive deterministic map $\{t_{x}\}_{x\geqslant 0}$, we write $T_{x}=O_{\mathbb{P}}(t_{x})$ if the collection $\{T_{x}/t_{x}\}_{x\geqslant 0}$ is tight, and $T_{x}=o_{\mathbb{P}}(t_{x})$ if $T_{x}/t_{x}\to 0$ in probability.
1.1.1 Spherically symmetric jumps
Denote the first coordinate of $\bm{\xi}$} by $\xi^{(1)}$, which is a real-valued random variable. We introduce the large deviation rate function
$I(x):=\sup_{\lambda>0}\Big(\lambda x-\log\phi_{\xi^{(1)}}(\lambda)\Big),$ (1)
where $\phi_{\xi^{(1)}}(\lambda):=\mathbb{E}[e^{\lambda{\xi^{(1)}}}]$ is the moment generating function for $\xi^{(1)}$}. In this case, we make the following assumptions on $\bm{\xi}$}:
  • (A3)
    the law of $\bm{\xi}$} is spherically symmetric in $\mathbb{R}^{d}$, and its radial component $R=|{\bm{\xi}}|$ satisfies $\mathbb{P}(R=0)<1$;
  • (A4)
    $\log\rho\in(\mathrm{ran}I)^{\circ}$, where $(\mathrm{ran}I)^{\circ}$ is the interior of the range of $I$. In other words, there exists $c_{1}>0$ such that $I(c_{1})=\log\rho$. It can be shown that $c_{1}\in(\mathrm{ran}(\log\phi_{\xi^{(1)}})^{\prime})^{\circ}$.
Note that under assumptions (A2) and (A3), $I$ is $C^{1}$ and convex in its domain, and attains a unique global minimum at x=0x=0.333See e.g., Theorem 26.3 of [rockafellar1970convex]. We will also see from Lemma 24 below that (A3) implies that both $\bm{\xi}$} and $\xi^{(1)}$} are non-lattice if $d\geqslant 2$. [^4] We say the law of $\bm{\xi}$} is non-lattice if there is no $\mathbf{y}\in\mathbb{R}^{d}$ such that the support of ${\bm{\xi}}+\mathbf{y}$ is contained in a discrete subgroup of $\mathbb{R}^{d}$. A consequence of (A4) is that $\rho>1$, and hence the particles form a supercritical Galton-Watson process. Denote by $c_{2}:=I^{\prime}(c_{1})$ the value of $\lambda$ at which the supremum in (1) is taken for $x=c_{1}$. In particular, (A4) implies that $\phi_{\xi^{(1)}}$ is well-defined in a neighborhood of $c_{2}$. The constants $c_{1}$ and $c_{2}$ here depend only on the offspring distribution (through $\rho$) and the increment distribution (through $\phi_{\xi^{(1)}}(\lambda)$) and will be fixed throughout this paper.
1.1.2 Non-spherically symmetric jumps
Denote by $\phi_{\bm{\xi}}(\bm{\lambda})=\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ the moment generating function for $\bm{\xi}$}. Let
$I(\mathbf{x}):=\sup_{\bm{\lambda}\in\mathbb{R}^{d}}\Big(\bm{\lambda}\cdot\mathbf{x}-\log\phi_{\bm{\xi}}(\bm{\lambda})\Big)=\sup_{\bm{\lambda}\in\mathbb{R}^{d}}\Big(\bm{\lambda}\cdot\mathbf{x}-\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]\Big)$ (2)
denote the large deviation rate function for $\bm{\xi}$}. We impose the following assumptions:
  • (A5)
    $\bm{\xi}$} satisfies the strong non-lattice condition $\limsup_{|\mathbf{t}|\to\infty}|\mathbb{E}[e^{i\mathbf{t}\cdot{\bm{\xi}}}]|<1$;
  • (A6)
    $\log\rho\in(\mathrm{ran}I(\cdot,\mathbf{0}))^{\circ}$, where $I(\cdot,\mathbf{0})$ refers to the function $I$ with the last d1d-1 variables fixed at zero; let $\widehat{c}_{1}>0$ satisfy $I(\widehat{c}_{1},\mathbf{0})=\log\rho$, then $(\widehat{c}_{1},\mathbf{0})\in(\mathrm{ran}\nabla\log\phi_{\bm{\xi}})^{\circ}$.
Denote by $\mathbf{c}_{2}=\nabla I(\widehat{c}_{1},\mathbf{0})$, which is the value of $\bm{\lambda}$ where the supremum (2) is attained at $\mathbf{x}=(\widehat{c}_{1},\mathbf{0})$. The intuition behind the vector $\mathbf{c}_{2}$ is explained by Figure 1 below. The assumption (A5) is crucial for the non-degeneracy of the jumps. More precisely, (A5) implies that the support of $\bm{\xi}$} cannot be contained in a proper subspace of $\mathbb{R}^{d}$. This is necessary for the BRW to be able to reach the target ball $B_{x}$ and shows that $d$ is the ‘real’ dimension of the increments. A consequence of (A6) is that $\phi_{\bm{\xi}}$ is well-defined in a neighborhood of $\mathbf{c}_{2}$. We will see in Proposition 21 that (A3) and (A4) together imply (A6) with $\widehat{c}_{1}=c_{1}$.
Figure 1: A two-dimensional visualization of the vector $\mathbf{c}_{2}$. The range of a (non-spherically symmetric) BRW roughly grows linearly in time (illustrated with the shaded ellipses); at time $\tau_{x}$, the ellipse is nearly tangent to the target ball $B_{x}$ (shaded disk). Roughly speaking, the vector $\mathbf{c}_{2}$ is normal to the tangent line.
1.2 Main results
In the following, we fix a dimension $d\geqslant 1$. Consider first the case of spherically symmetric jumps. Let us define
$A(x):=\frac{x}{c_{1}}+\frac{d+2}{2c_{2}c_{1}}\,\log x.$ (3)
For a finite set $G$, we use #G\#G to denote its cardinality. Denote by $S:=\{\forall n\geqslant 1,\penalty 10000\ \#V_{n}>0\}$ the survival event that the underlying branching process survives at all times. Since $\rho>1$, we have $p:=\mathbb{P}(S)>0$.
Theorem 1.
Assume (A1)–(A4). Conditional upon survival, it holds that
$\tau_{x}=A(x)+O_{\mathbb{P}}(\log\log x)=\frac{x}{c_{1}}+\frac{d+2}{2c_{2}c_{1}}\,\log x+O_{\mathbb{P}}(\log\log x).$ (4)
In other words, the collection $\{(\tau_{x}-A(x))/\log\log x\}_{x>0}$ is tight.
A natural question is whether the $O_{\mathbb{P}}(\log\log x)$ can be replaced by a $O_{\mathbb{P}}(1)$, which we later formulate as Conjecture 1.555We remark that the conjecture has been resolved by [blanchet2024tightness] after the first version of this paper was posted; see Remark 4. We discuss a few pieces of evidence that support this conjecture, and further aspects will be discussed in Section 5. First, this is the case for $d=1$ (Proposition 7 below) and for the branching Brownian motion in $\mathbb{R}^{d}$ (Theorem 1 of [zhang2024modeling]). A second piece of evidence is the following tightness result on first passage times that is also central to our proof of Theorem 1.
Theorem 2.
For $r\in(0,1)$ and $x>0$, let $t^{(r)}_{x}$ be the $r$-th quantile of the first passage time $\tau_{x}$, i.e., $t^{(r)}_{x}=\inf\{t:\mathbb{P}(\tau_{x}\leqslant t\mid S)\geqslant r\}$. Assume (A1), (A2), and (A6). There exist constants $C,c>0$ independent of $x$ such that for each $y\in[0,x]$,
$\mathbb{P}\left(|\tau_{x}-t^{(1/2)}_{x}|>y\mid S\right)\leqslant Ce^{-cy}.$ (5)
In particular, conditioned upon survival, the collection of laws $\{\tau_{x}-t^{(r)}_{x}\}_{x>0}$ is tight for each $r\in(0,1)$.
The analog of Theorem 2 for the maximum of BRW traces back to [dekking1991limit] and [mcdiarmid1995minimal], which assumed uniformly bounded jumps. The case of unbounded jumps was later resolved by [bramson2007tightness] and [bramson2009tightness] using a different approach. Theorem 2 generalizes these results to the first passage times of BRW in arbitrary dimensions. In this setting, for the jumps, we only require mean zero and assumption (A6). Since (A3) and (A4) together imply (A6), Theorem 2 applies in the spherically symmetric case as well.
Extending these results to the case of non-spherically symmetric jumps, we define
$\widehat{A}(x):=\frac{x}{\widehat{c}_{1}}+\frac{d+2}{2\,\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\,\log x,$ (6)
where we recall the definition of $\widehat{c}_{1}$ from assumption (A6).
Theorem 3.
Assume (A1), (A2), (A5), and (A6). Conditional upon survival,
$\tau_{x}=\widehat{A}(x)+o_{\mathbb{P}}(\log x)=\frac{x}{\widehat{c}_{1}}+\frac{d+2}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\,\log x+o_{\mathbb{P}}(\log x).$ (7)
In other words, the collection $(\tau_{x}-\widehat{A}(x))/\log x$ converges to 0 in probability as $x\to\infty$.
There are certain instances for non-spherically symmetric jumps where the asymptotic (7) can be improved. Such a class includes product measures (meaning that $\bm{\xi}$} has independent coordinates) and non-degenerate centered elliptical distributions on $\mathbb{R}^{d}$. [^6] We say that a centered random variable $\bm{\xi}$} follows a non-degenerate elliptical distribution if ${\bm{\xi}}\overset{\mathrm{law}}{=}TU$ where $U$ is spherically symmetric and $T$ is an invertible linear transformation. These will be addressed in detail in Section 3.2.
An immediate corollary of Theorems 2 and 3 is the following strong law of large numbers for the first passage times.
Corollary 4.
Assume (A1), (A2), (A5), and (A6). Conditioned upon survival,
$\frac{\tau_{x}}{x}\to\frac{1}{\widehat{c}_{1}}\qquad\text{a.s.}$
As yet another consequence of Theorems 2 and 3 (while slightly modifying their proofs), the full range of the BRW forms a dense subset of $\mathbb{R}^{d}$, complementing Corollary 2.5 of [oz2023] that the full range of the BBM is dense.
Corollary 5.
Assume (A1), (A2), (A5), and (A6), and that the MGF $\phi_{{\bm{\xi}}}$ is well-defined in a neighborhood of $\mathbf{0}$. Conditioned upon survival, the set
$R:=\bigcup_{n\geqslant 0}\bigcup_{v\in V_{n}}\{\bm{\eta}_{v,n}(n)\}$
is dense in $\mathbb{R}^{d}$ almost surely.
In Section 4, we test Theorems 1 and 3 numerically with a path purging algorithm that significantly improves the computational efficiency for capturing the extremal behavior of a BRW.
1.3 Proof strategy
In this subsection, we discuss the high-level intuition of the strategy we will follow, building from the case of $d=1$. Throughout, we condition on the survival event. We first assume that the jump distribution is spherically symmetric. For one-dimensional BRW, we expect the following parity relation between maximum and FPT. Recall that the asymptotic for the maximum $M_{n}$ at time $n$ is
$M_{n}=m_{n}+O_{\mathbb{P}}(1),\quad\text{ where }\quad m_{n}:=c_{1}n-\frac{3}{2c_{2}}\,\log n.$ (8)
Inverting (8) leads to the asymptotic (e.g., using the Lambert $W$ function)
$n=\frac{M_{n}}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log M_{n}+O_{\mathbb{P}}(1).$
It is then natural to guess that for one-dimensional BRW,
$\tau_{x}=t_{1}(x)+O_{\mathbb{P}}(1),\quad\text{ where }\quad t_{1}(x):=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x,$ (9)
which we will prove in Proposition 7 below in Section 2.1. Recall (3). Let us decompose $A(x)=t_{1}(x)+t_{2}(x)$, where
$t_{2}(x):=A(x)-t_{1}(x)=\frac{d-1}{2c_{2}c_{1}}\,\log x.$ (10)
For a general dimension $d\geqslant 1$, we shall explain the extra term $t_{2}(x)$.
The main difficulty in analyzing the first passage times in higher dimensions is that the second moment method does not immediately apply if one counts the number of particles in the target ball $B_{x}$ at a certain time. To bypass this difficulty, we prove separately the upper and lower bounds for $\tau_{x}$.
A key step in these bounds is Proposition 8, which keeps track of the number of particles that are near the frontier (say, a distance of $\log n$ away) for a one-dimensional BRW. Roughly speaking, there are around $xe^{c_{2}x}$ particles that are of distance $x$ behind the frontier, where $x\ll\log n$. This amounts to generalizing the proof of the asymptotic of the maximum (8); see [bramson2016convergence]. A notable distinction is that unfortunately, one cannot conclude from the second moment method the lower bound for the number of particles near the frontier with overwhelming probability. Instead, one only concludes the lower bound with a uniformly positive probability. This gap can be filled using Theorem 2, which shows that a uniformly positive probability would already be sufficient.
A particularly nice feature of the spherical symmetry of the jumps is that the displacements in different directions are approximately independent (likewise, the same result Theorem 1 holds if we instead assume independent centered jumps in the $d$ directions). In particular, at time $n$, one roughly expects a proportion $O(n^{-(d-1)/2})$ of the particles that lie within unit distance from the origin in the last d1d-1 dimensions. In other words, one expects that the first passage event is realized around the time when $x^{(d-1)/2}$ particles reach a distance of $x$ for the $\mathbb{R}^{d}$-valued BRW projected onto the first dimension; see Figure 2 below. On the other hand, by (8), the extra time $t_{2}(x)$ from (10) pushes the frontier ahead by a distance of approximately $c_{1}t_{2}(x)=(d-1)(\log x)/(2c_{2})$, which amounts to having approximately $x^{(d-1)/2}$ extra particles in view of Proposition 8.
Figure 2: Schematic description of the heuristic for spherically symmetric jumps: around a proportion $x^{-(d-1)/2}$ of the particles that reach roughly $x$ far in the first coordinate (represented by blue dots) lie in the ball $B_{x}$. The path in red represents (part of) the trajectory that realizes the FPT, whereas the other branches are shown in black.
If the jump distribution is not spherically symmetric, applying similar arguments above (with Proposition 8 replaced by Theorem B of [biggins1979growth] and Proposition 10 by Proposition 17; we omit the details here) will lead to
$\tau_{x}=\frac{x}{\widehat{c}_{1}}+O_{\mathbb{P}}(\log x),$ (11)
where the corresponding upper and lower bounds for $\tau_{x}$ are separated by an order of $\log x$. To pin down the coefficient of the logarithmic correction term, we extend the arguments for the maximum of one-dimensional BRW ([bramson2016convergence]) with a modified barrier. The modified barrier event lies in $\mathbb{R}^{d}$ and hence results on random walks in cones ([denisov2015random]) are applicable. Note that the same proof also applies for the spherically symmetric case, but would yield less precise asymptotics compared to (4) (an error of $o_{\mathbb{P}}(\log x)$ instead of $O_{\mathbb{P}}(\log\log x)$).
1.4 Applications to polymer physics
Our interest in estimating the FPT of spatial branching processes stems from their relevance to polymer physics. The mechanical behavior of polymer networks appears to be controlled by the shortest paths between distant nodes in the polymer network, as shown by recent works [yin2020topological, yin2024network] using coarse-grained molecular dynamics (CGMD) simulations. The shortest path length (SP) serves as an indicator of how stretched the average load-bearing chains are in the polymer network. Theories based on SP go beyond traditional polymer models focused on elastic behaviors and aim to explain the inelastic behaviors exhibited by such materials.
The SP between (fixed) distant nodes is equivalent to the FPT of the BRW beginning from the origin (source node) and terminating at a desired point in space (destination node). Moreover, the occurrence of the cross-links along the polymer backbone mimics a Poisson process. As a result, a cross-link between two chains is equivalent to a branching event observed in the BRW genealogy tree. The agreement of the mean FPT from the BRW theory with the mean SP from the CGMD simulation results has been shown in our recent work [zhang2024modeling]. Furthermore, the numerical implementation of the BRW model shows an excellent agreement between its FPT distribution and the SP distribution from the CGMD simulation. Our earlier work initially focused on constant length jumps with the BRW model where there existed a correlation between the components of a jump step in the $d$ dimensions by virtue of being sampled from the unit sphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\left\lVert\mathbf{x}\right\rVert=1\}$. This was later shown to be represented equally well by the Gaussian BRW and BBM in which the components of each jump step are independent in the $d$ dimensions. However, the analysis was restricted to spherically symmetric jumps to accurately model the as-prepared polymer network prior to deformation.
2 Proofs of main results
This section is devoted to the proofs of the main results. We will first analyze the simple case of $d=1$ in Section 2.1, where a precise asymptotic of the FPT follows directly from inverting the maxima asymptotic (8). We then collect the necessary tools in Sections 2.2.1 and 2.2.2, and prove Theorems 1 and 2 in Sections 2.2.3 and 2.3 respectively. Most of the arguments before and including Section 2.2.1 closely follow the approach of [bramson2016convergence]. [^7] The results in [bramson2016convergence] are stated for the case $p_{0}=0$, but the general case follows in a similar way. The case of non-spherically symmetric jumps (Theorem 3) will be addressed in Section 2.4, where we also summarize preliminary results on random walks that are constrained to stay in a cone. Section 2.5 presents the proofs of Corollaries 4 and 5.
We begin by introducing some necessary notations. Following [bramson2016convergence], we denote by $V_{n}$ the offspring on the $n$-th generation. Let $\{\bm{\eta}_{v,n}(k)\}_{k=0,\dots,n}$ denote the $d$-dimensional random walk where $\bm{\eta}_{v,n}(k)$ is the position of the $k$-th generation for the path in the tree leading to a vertex $v\in V_{n}$. We also let $\eta^{(1)}_{v,n}(k)$ denote the first coordinate of $\bm{\eta}_{v,n}(k)$. When $d=1$, these notions coincide. For deterministic quantities or functions $A,B$, We use Vinogradov’s symbol $A\ll B$ (or $A=O(B)$) to denote $|A|\leqslant CB$ with some constant $C>0$ that depends only on the distribution of $\bm{\xi}$} and the dimension $d$. We write $A\asymp B$ if $A\ll B\ll A$. The notation $\left\lVert\cdot\right\rVert$ always denotes the Euclidean norm. We use $B_{\mathbf{x}}(r)$ to denote the shifted ball of radius $r$ centered at $\mathbf{x}\in\mathbb{R}^{d}$, $B_{\mathbf{x}}=B_{\mathbf{x}}(1)$, and $B_{x}=B_{(x,\mathbf{0})}$ for $x\in\mathbb{R}$. From here on, without loss of generality, we will assume that quantities that tend to infinity are integers. The arguments can be made rigorous by writing floor functions everywhere, but we have chosen not to do so for the simplicity of the notation.
2.1 Warming up in dimension one
For $\beta\geqslant 1$ and $n\in\mathbb{N}$ we define the event
$G_{n,\beta}:=\bigcup_{v\in V_{n}}\bigcup_{0\leqslant k\leqslant n}\left\{\eta^{(1)}_{v,n}(k)\geqslant\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}.$
Lemma 6 (Lemma 2.4 of [bramson2016convergence]).
Assume conditions (A1), (A2), and (A4), and assume that the jump distribution is non-lattice. It holds that for all $n\in\mathbb{N}$, $\mathbb{P}(G_{n,\beta})\ll\beta e^{-c_{2}\beta}$ for $\beta\geqslant 1$.
Proposition 7.
Suppose that a one-dimensional BRW satisfies conditions (A1), (A2), and (A4), $p_{0}=0$ (no extinction) and that the jump distribution is non-lattice. We have
$\tau_{x}=t_{1}(x)+O_{\mathbb{P}}(1)=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x+O_{\mathbb{P}}(1).$
Proof.
For the lower bound, we fix $\varepsilon>0$ and pick $\beta$ large enough such that $\mathbb{P}(G_{n,\beta})<\varepsilon$ for all $n\in\mathbb{N}$, by means of Lemma 6. In other words, with a probability larger than $1-\varepsilon$, the whole BRW stays below the curve
$k\mapsto\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}.$
It is clear that with $C_{1}=C_{1}(\beta)>0$ chosen large enough,
$\max_{1\leqslant k\leqslant t_{1}(x)-C_{1}}\Big(\frac{km_{t_{1}(x)-C_{1}}}{t_{1}(x)-C_{1}}+\beta+\frac{4}{c_{2}}(\log(\min(k,t_{1}(x)-C_{1}-k)))_{+}\Big)\leqslant x.$
This yields $\mathbb{P}(\tau_{x}\leqslant t_{1}(x)-C_{1})\leqslant\mathbb{P}(G_{n,\beta})<\varepsilon$, proving the lower bound.
For the upper bound, we fix $\varepsilon>0$. Pick $C_{2}>0$ such that $\mathbb{P}(M_{t_{1}(x)+C_{2}}>x)>1-\varepsilon/3$ uniformly in $x$ by (8) and the convergence in law of the recentered maximum (Theorem 1.1 of [bramson2016convergence]). Using the same analysis for the lower bound, there exists $C_{3}>0$ such that $\mathbb{P}(M_{t_{1}(x)+C_{2}}>x+C_{3})<\varepsilon/3$. Therefore, with probability $1-2\varepsilon/3$, there is $v\in V_{t_{1}(x)+C_{2}}$ such that $x\leqslant\eta^{(1)}_{v,t_{1}(x)+C_{2}}(t_{1}(x)+C_{2})\leqslant x+C_{3}$. Since the distribution of $\xi$} is non-lattice, by the Chung–Fuchs theorem (Theorem 5.4.8 of [durrett2019probability]), the random walk with step distribution $\xi$ initiated from $v$ is recurrent, and hence there exists $C_{4}>0$ such that the descendants of $v$ reaches [x1/2,x+1/2][x-1/2,x+1/2] with probability $1-\varepsilon/3$ within time $C_{4}$. Combining the above leads to $\mathbb{P}(\tau_{x}>t_{1}(x)+C_{2}+C_{4})<\varepsilon$, as desired. ∎
2.2 Proof of Theorem 1
We now present the proof of Theorem 1 for $d\geqslant 2$, provided that Theorem 2 holds true. The proof of Theorem 2 will be the focus of the next section. The following auxiliary results are necessary as explained in our proof strategy.
2.2.1 The number of particles near maximum
The goal of this section is to prove the following result on the number of particles near the frontier of a one-dimensional BRW, which might be of independent interest.
Proposition 8.
Assume (A1)–(A4). There exists $L>1$ depending only on the law of the BRW such that the following holds conditioned upon survival. Given any $\varepsilon>0$, there exists $C>0$ independent from $n$ and $x$ such that uniformly for $n$ large enough and for $x\in[2,\sqrt{n}]$,
$\mathbb{P}\left(\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>Cxe^{c_{2}x}\mid S\right)<\varepsilon$ (12)
and
$\mathbb{P}\left(\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>\frac{1}{C}xe^{c_{2}x}\mid S\right)>\frac{1}{L}.$ (13)
Remark 1.
Proposition 8 may be compared against [biggins1979growth, rouault1993precise], where it is proven that for a location that is $\Omega(n)$ away from $m_{n}$, the number of particles therein has the same order as its expectation. Proposition 8 shows that this is not the case for locations that are $O(\sqrt{n})$ near $m_{n}$. On the other hand, the number of particles that are $O(\sqrt{n})$ away from the origin can be described by the BRW central limit theorem [biggins1990central]. An interesting question would be to analyze the phase transitions between those regimes. We also expect that (13) holds if we replace its right-hand side by $1-\varepsilon$, but the current version suffices for our purpose.
In the rest of this section, we assume that $\mathbb{P}$ is the probability measure conditioned upon the survival event $S$. Following the setup of Section 2.2 of [bramson2016convergence], we introduce $\lambda_{n}$ which is the value of $\lambda$ where the supremum of (1) is attained with $x=m_{n}/n$. By (15) of [bramson2016convergence], it holds $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. For a given $v\in V_{n}$, let $\mathbb{Q}^{(n)}$ be a probability measure on length-$n$ random walk paths defined by
$\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}^{(n)}}:=e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})-nI(m_{n}/n)}\asymp n^{3/2}\rho^{-n}e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})}.$ (14)
It follows that under $\mathbb{Q}^{(n)}$, $\{\eta^{(1)}_{v,n}(k)-km_{n}/n\}_{k=0,\dots,n}$ is a mean zero random walk. Let
$Q_{n,\beta}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\penalty 10000\ \eta^{(1)}_{v,n}(k)<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}$
and define
$g(n,\beta,x):={\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]}.$ (15)
We first reduce the proof of the upper bound (12) into proving the following estimate on $g(n,\beta,x)$.
Proposition 9.
Assume (A1)–(A4). Uniformly in $x\in[2,\sqrt{n}]$,
$g(n,\beta,x)\ll\beta(x+\beta)e^{c_{2}x}.$
Proof of (12) given Proposition 9.
For fixed $C,\varepsilon>0$, we pick $\beta$ large enough so that $\mathbb{P}(G_{n,\beta})<\varepsilon/2$ by Lemma 6. On the event $G_{n,\beta}^{c}$, $Q_{n,\beta}=V_{n}$. By Proposition 9, uniformly for $n$ and $x\in[2,\sqrt{n}]$,
$\mathbb{E}\left[\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}{\mathbbm{1}}_{G_{n,\beta}^{c}}\right]\ll\beta(x+\beta)e^{c_{2}x}.$
Markov’s inequality then yields that for some $L>0$,
$\mathbb{P}\left(G_{n,\beta}^{c}\cap\left\{\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>L\beta(x+\beta)e^{c_{2}x}\right\}\right)<\frac{\varepsilon}{2}.$
This yields (12). ∎
Proof of Proposition 9.
Let $\{S_{k}\}_{1\leqslant k\leqslant n}$ be a random walk with increment $\xi^{(1)}$. By the many-to-one formula, we have
$\hskip 14.22636pt\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]$
$=\rho^{n}\mathbb{P}\Big(S_{k}<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\text{ for }1\leqslant k\leqslant n;\,S_{n}\geqslant m_{n}-x\Big)$
$=\sum_{j=0}^{x+\beta-1}\rho^{n}\mathbb{P}\Big(S_{k}<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\text{ for }1\leqslant k\leqslant n;$
$\hskip 113.81102ptS_{n}\in[m_{n}+\beta-j-1,m_{n}+\beta-j)\Big).$
By substituting $i=\beta-j$, the latter probability is precisely $\chi_{n,n}^{\mathbb{P}}(i)$ defined below (15) in [bramson2016convergence]. It is proved in (18) therein the estimate
$\chi_{n,n}^{\mathbb{P}}(i)\ll\beta(\beta-i+2)\rho^{-n}e^{-c_{2}i},$
where our assumption $x\leqslant\sqrt{n}$ is used. This leads to
$\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]\ll\sum_{j=0}^{x+\beta}\beta(j+2)e^{-c_{2}(\beta-j)}\ll\beta(x+\beta)e^{c_{2}x},$
completing the proof. ∎
Proof of (13).
We apply the second moment method. Define
$P_{n}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\penalty 10000\ \eta^{(1)}_{v,n}(k)<\frac{km_{n}}{n}\right\}.$
For $x\in[2,\sqrt{n}]$ and $v\in V_{n}$, define the event
$H_{v,n}(x):=\{v\in P_{n},\penalty 10000\ \eta^{(1)}_{v,n}(n)\in[m_{n}-x,m_{n}-x+1)\}$
and $\Delta_{n,x}:=\sum_{v\in V_{n}}{\mathbbm{1}}_{H_{v,n}(x)}$. It follows that
$\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}\geqslant\Delta_{n,x}.$ (16)
Let us first compute the first moment $\mathbb{E}[\Delta_{n,x}]$. Using (14), we have
$\mathbb{P}(H_{v,n}(x))$ $\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}^{(n)}}\left[e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})}{\mathbbm{1}}_{H_{v,n}(x)}\right]$
$\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}^{(n)}}\left[e^{c_{2}x}{\mathbbm{1}}_{H_{v,n}(x)}\right]=n^{3/2}\rho^{-n}e^{c_{2}x}\mathbb{Q}^{(n)}(H_{v,n}(x)).$
By the ballot theorem, in the form of Lemma 2.1 of [bramson2016convergence], We have
$\mathbb{Q}^{(n)}(H_{v,n}(x))\gg n^{-3/2}\max(x-1,1)\gg n^{-3/2}x.$
Combining the above yields
$\mathbb{E}[\Delta_{n,x}]=\rho^{n}\mathbb{P}(H_{v,n}(x))\gg xe^{c_{2}x}.$ (17)
Next we estimate the second moment $\mathbb{E}[\Delta_{n,x}^{2}]$. Observe that by (A1) (see (29) of [bramson2016convergence]),
$\mathbb{E}[\Delta_{n,x}^{2}]\ll\mathbb{E}[\Delta_{n,x}]+\rho^{n}\sum_{s=1}^{n}\rho^{s}\mathbb{P}(H_{v,n}(x)\cap H_{w,n}(x)\text{ for }v\sim_{s}w),$ (18)
where $v\sim_{s}w$ means that the distance of $v$ and $w$ in the genealogical tree is equal to $2s$. To bound the latter probability we condition on the location of the common ancestor of $v$ and $w$, $\eta^{(1)}_{v,n}(n-s)$. We have $\text{for }v\sim_{s}w$,
$\hskip 14.22636pt\mathbb{P}(H_{v,n}(x)\cap H_{w,n}(x))$
$\leqslant\sum_{j=0}^{\infty}\mathbb{P}\left(\eta^{(1)}_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\penalty 10000\ \eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
$\hskip 113.81102pt\times\left(\sup_{y\in[-(j+1),-j)}P(n,s,x,y)\right)^{2},$
where
$P(n,s,x,y)$ $:=\mathbb{P}\Bigg(\eta^{(1)}_{v,n}(n-s+\ell)\leqslant\frac{(n-s+\ell)m_{n}}{n}\text{ for }0\leqslant\ell\leqslant s,$
$\hskip 34.14322pt\eta^{(1)}_{v,n}(n)\in[m_{n}-(x+1),m_{n}-x)\mid\eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}=y\Bigg).$
With the same change of measure and ballot theorem argument as above, we get
$\hskip 14.22636pt\mathbb{P}\left(\eta^{(1)}_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\penalty 10000\ \eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
$\ll e^{3(n-s)(\log n)/(2n)}\rho^{-(n-s)}e^{\lambda_{n}j}$
$\hskip 56.9055pt\times(\max(1,n-s))^{-3/2}\max\left((m_{n-s}-(n-s)m_{n}/n+j),1\right)$
$\ll\rho^{-(n-s)}e^{\lambda_{n}((3/(2c_{2}))((n-s)(\log n)/n-\log(n-s))+j)}$
$\hskip 56.9055pt\times\max\left(\big(j+\frac{(n-s)\log n}{n}-\log(n-s)\big),1\right)$
$\ll\rho^{-(n-s)}je^{c_{2}j-3\log(n-s)/2+3(n-s)\log n/(2n)},$
where in the last step we use $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. In a similar manner, (while using (6) of [bramson2016convergence]), we obtain the estimate
$P(n,s,x,y)\ll\rho^{-s}jxe^{c_{2}(x-j)+3((s\log n)/n-\log s)/2}.$
In conclusion, we arrive at
$\mathbb{E}[\Delta_{n,x}^{2}]$ $\ll xe^{c_{2}x}+\sum_{s=1}^{n}\sum_{j=0}^{\infty}je^{c_{2}j-3\log(n-s)/2+3(n-s)\log n/(2n)}$
$\hskip 113.81102pt\times\left(jxe^{c_{2}(x-j)+3((s\log n)/n-\log s)/2}\right)^{2}$
$\ll xe^{c_{2}x}+x^{2}e^{2c_{2}x}\sum_{s=1}^{n}\frac{n^{3/2}e^{(3s\log n)/(2n)}}{(\max(1,n-s))^{3/2}s^{3}}\ll x^{2}e^{2c_{2}x},$ (19)
where in the last step we argue similarly as in the proof of Lemma 2.7 of [bramson2016convergence]. Combined with (16), (17), and the Paley-Zygmund inequality leads to (13). ∎
2.2.2 Large deviation computations
In the following, we let $\{(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$}, and $(F_{n},\mathbf{J}_{n})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Recall the definition of $m_{n}$ from (8). We may slightly abuse notation and use $B_{\mathbf{x}}(r)$ to denote the ball centered at $\mathbf{x}\in\mathbb{R}^{d-1}$ with radius $r$ in dimension d1d-1 (instead of dimension $d$).
Proposition 10.
Fix a large constant $C>0$ and a dimension $d\geqslant 1$. Suppose that the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$} satisfies conditions (A2)–(A4) from Section 1.1.
  1. (i)
    Uniformly for any positive sequence $u(n)\leqslant C\sqrt{n}$ and any $c=c(n)\in[-C\log n,C\log n]$,
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\mid F_{n}\geqslant m_{n}+c)\ll u(n)^{d-1}n^{-(d-1)/2}.$ (20)
  2. (ii)
    Fix $a(n)\gg 1$ and $a(n)=o(\log n)$. Uniformly for $\mathbf{y}$ with $\left\lVert\mathbf{y}\right\rVert\leqslant C\sqrt{n}$ and $c\in[-C\log n,C\log n]$,
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1)\mid F_{n}\in[m_{n}+c,m_{n}+c+a(n)])\gg n^{-(d-1)/2}.$ (21)
  3. (iii)
    For any $\varepsilon>0$, there exists $K>0$ such that uniformly for $c\in[-C\log n,C\log n]$,
    $\mathbb{P}(\mathbf{J}_{n}\not\in B_{\mathbf{0}}(K\sqrt{n})\mid F_{n}\geqslant m_{n}+c)<\varepsilon.$ (22)
The proof of Proposition 10 is deferred to the appendix.
2.2.3 Proof of Theorem 1 conditioned on Theorem 2
Recall (3). Let us define
$A_{1}(x):=A(x)+\frac{2}{c_{2}c_{1}}\,\log\log x,\quad A_{2}(x):=A(x)-\frac{1}{c_{2}c_{1}}\,\log\log x.$
In view of Theorem 2, it suffices to prove that there exists $L>0$ such that for any $a(x)$ that tends to infinity, and any $\varepsilon>0$, there exists $N>0$ such that for any $x\geqslant N$,
$\mathbb{P}(\tau_{x}>A_{1}(x)+a(x))\leqslant 1-\frac{1}{L}$ (23)
and
$\mathbb{P}(\tau_{x}<A_{2}(x)-a(x))<\varepsilon.$ (24)
In the following, we fix $\varepsilon>0$ and $a(x)$.
Proof of the upper bound. Define
$\alpha(x):=\frac{d-1}{2c_{2}}\,\log x-\frac{1}{c_{2}}\,\log\log x+c_{1}a(x)$
and
$W_{n}:=\left\{v\in V_{n}:\bm{\eta}_{v,n}(n)\in\Big(m_{n}-\alpha(x),m_{n}\Big)\times B_{\mathbf{0}}(\sqrt{x})\right\}.$
We first prove that for $x$ large enough and some large constant $L_{1}$ to be determined,
$\mathbb{P}\big(\#W_{(\log x)^{2}}<L_{1}x^{({d-1})/{2}}\big)<1-\frac{1}{L},$ (25)
for some constant $L>0$ depending on $L_{1}$ only. Indeed, let
$\widetilde{W}_{n}:=\left\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\in\Big(m_{n}-\alpha(x),m_{n}\Big)\right\}.$
It follows from Proposition 8 that for $x$ large enough, with $n=(\log x)^{2}$, $\mathbb{P}(\#\widetilde{W}_{n}<L_{1}x^{(d-1)/2})<1-1/L$. Comparing the definitions of $W_{n}$ and $\widetilde{W}_{n}$, it remains to show that at generation $n=(\log x)^{2}$, the maximum norm of the BRW is smaller than $\sqrt{x}$ with probability 1o(1)1-o(1). But this is a direct consequence of Theorem 1.1 of [zhang2024large].
The plan is to evolve (on the event of (25)) the $L_{1}x^{(d-1)/2}$ particles in $W_{(\log x)^{2}}$ (independently) so that at least one of those families has a descendant landing in $B_{x}$ at time $A_{1}(x)+a(x)$ with high probability, where we recall that $B_{x}$ is the ball of radius one centered at $(x,\mathbf{0})$. Let us now fix $w\in W_{(\log x)^{2}}$. Define $\tau^{\prime}(x):=A_{1}(x)+a(x)-(\log x)^{2}+K_{1}$ for some large constant $K_{1}>0$ (to be determined) and
$p(w):=\mathbb{P}\left(\bm{\eta}_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{\tau^{\prime}(x)}\right),$
which is the probability that at least one descendant of $w$ lands in $B_{x}$ at time $A_{1}(x)+a(x)+K_{1}$, or equivalently, the probability that a BRW reaching the set $B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})$ at generation $\tau^{\prime}(x)$. By independence and by picking the constant $L_{1}$ in (25) large enough, it suffices to prove $p(w)\gg x^{-(d-1)/2}$ for $w\in W_{(\log x)^{2}}$. Note that $\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})\in(m_{(\log x)^{2}}-\alpha(x),m_{(\log x)^{2}})$ and that for $K_{1}$ chosen large enough,
$m_{\tau^{\prime}(x)}$ $=x+\frac{d-1}{2c_{2}}\log x+\frac{2}{c_{2}}\log\log x-c_{1}(\log x)^{2}+c_{1}a(x)+c_{1}K_{1}+\frac{3\log c_{1}}{2c_{2}}+o(1)$
$\geqslant x+\frac{d-1}{2c_{2}}\log x+\frac{2}{c_{2}}\log\log x-c_{1}(\log x)^{2}+c_{1}a(x)+1$
$=x-(m_{(\log x)^{2}}-\alpha(x))+1.$
Therefore, the asymptotic of the maximum at time $\tau^{\prime}(x)$, $m_{\tau^{\prime}(x)}$, is at least $x+1-\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})$. Using the lower bound on $\Delta_{n,x}$ derived in the proof of (13), there is some constant $L>0$ such that
$\mathbb{P}\left(\eta^{(1)}_{v,\tau^{\prime}(x)}(\tau^{\prime}(x))\in\big(x-\frac{1}{2},x+\frac{1}{2}\big)-\eta^{(1)}_{w,(\log x)^{2}}((\log x)^{2})\text{ for some }v\in V_{\tau^{\prime}(x)}\right)\geqslant\frac{1}{L}.$
Conditioned on the above event, Proposition 10(ii) shows that the particle $v$ that realizes this event has a chance of $\gg x^{-(d-1)/2}$ to land in $B_{x}-\bm{\eta}_{w,(\log x)^{2}}((\log x)^{2})$ at time $\tau^{\prime}(x)$, leading to $p(w)\geqslant x^{-(d-1)/2}/L$. This establishes (23).
Proof of the lower bound. Our goal is to bound $\mathbb{P}(\tau_{x}\leqslant A_{2}(x)-a(x))$ from above. According to Propositions 8 and 10, at time $A_{2}(x)$ there will be less than $O(1)$ number of particles that lie in $x,\infty)\times B_{\mathbf{0}}(1)$ in expectation. We need to argue that other particles that stayed in $B_{x}$ earlier but currently do not stay in $x,\infty)\times B_{\mathbf{0}}(1)$ do not contribute asymptotically. This is because most of the $O(1)$ particles will arrive not much earlier than $t_{1}(x)+t_{2}(x)$, and will not move quickly.
Consider $n=A_{2}(x)$. Recall that $\varepsilon>0$ and $a(x)\to\infty$ are fixed. We define $T_{v,n}(x):=\min\{0\leqslant k\leqslant n:\eta^{(1)}_{v,n}(k)\geqslant x\}$ and for an integer $j\in[a(x),t_{2}(x)]$,
$V_{n,j}:=\big\{v\in V_{n}:T_{v,n}(x)=A_{2}(x)-j\big\}.$
It follows from Lemma 6 that for any $\varepsilon>0$, there exists $L_{2}$ large enough such that
$\hskip 14.22636pt\mathbb{P}(\tau_{x}\leqslant A_{2}(x)-a(x))$
$\leqslant\mathbb{P}\left(\bigcup_{j=a(x)}^{t_{2}(x)+L_{2}}\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\mathbb{P}\left(M_{A_{2}(x)-t_{2}(x)-L_{2}}\geqslant x\right)$
$\leqslant\sum_{j=a(x)}^{t_{2}(x)+L_{2}}\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x}\}\right)+\frac{\varepsilon}{4}.$ (26)
Let us further decompose the event $\{\bm{\eta}_{v,n}(k)\in B_{x}\}$ depending on the location at the first time the random walk $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ reaches $x,\infty)\times\mathbb{R}^{d-1}$, whether $\bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)$ with some $c_{3}>0$ to be determined.
First, with $j$ fixed,
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\penalty 10000\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
$\leqslant\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\{\bm{\eta}_{v,n}(A_{2}(x)-j)\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
$=\mathbb{P}\left(\bigcup_{v\in V_{A_{2}(x)-j}}\{\bm{\eta}_{v,A_{2}(x)-j}(A_{2}(x)-j)\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right).$
Using Proposition 8 (while adjusting the proof suitably), we see that for some $L_{3}>0$,
$\mathbb{P}\left(\#\{v\in V_{A_{2}(x)-j}:\eta^{(1)}_{v,A_{2}(x)-j}(A_{2}(x)-j)\geqslant x\}\geqslant L_{3}j^{2}e^{-c_{2}c_{1}j}n^{(d-1)/2}\right)<\frac{\varepsilon}{10j^{2}}.$
Applying Proposition 10(iii) and the union bound then yields
$\mathbb{P}\left(\bigcup_{v\in V_{A_{2}(x)-j}}\{\eta^{(1)}_{v,A_{2}(x)-j}(A_{2}(x)-j)\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)\}\right)$
$\hskip 170.71652pt\leqslant\frac{\varepsilon}{10j^{2}}+L_{3}j^{2}e^{-c_{2}c_{1}j}(c_{3}j)^{d-1}.$
Second, using that $j\geqslant a(x)$ and $a(x)\to\infty$ as $x\to\infty$, as well as Theorem 3.2 of [gantert2018large], we have for some $c_{3},c_{4}>0$ large and independent of $x$, such that for all $x$ large enough,
$\hskip 14.22636pt\mathbb{P}\left(\bigcup_{v\in V_{n,j}}\bigcup_{T_{v,n}(x)\leqslant k\leqslant n}\{\bm{\eta}_{v,n}(k)\in B_{x},\penalty 10000\ \bm{\eta}_{v,n}(T_{v,n}(x))\in[x,\infty)\times B_{\mathbf{0}}(c_{3}j)^{c}\}\right)$
$\leqslant\mathbb{P}\left(\max_{v\in V_{j}}\left\lVert\bm{\eta}_{v,j}(j)\right\rVert\geqslant(c_{3}-1)j\right)$
$\leqslant d\,\mathbb{P}\left(\max_{v\in V_{j}}|\eta^{(1)}_{v,j}(j)|\geqslant\frac{(c_{3}-1)j}{\sqrt{d}}\right)$
$\leqslant e^{-j/c_{4}}.$
Inserting the above estimates to (26) leads to
$\mathbb{P}(\tau_{x}\leqslant A_{2}(x)-a(x))$ $\leqslant\frac{\varepsilon}{4}+\sum_{j=a(x)}^{\infty}\left(\frac{\varepsilon}{10j^{2}}+L_{3}j^{2}e^{-c_{2}c_{1}j}(c_{3}j)^{d-1}+e^{-j/c_{4}}\right).$
For any $\varepsilon>0$ and $a(x)\to\infty$, the right-hand side is $\varepsilon$ for $x$ large enough. This completes the proof of the lower bound (24).
2.3 Tightness of first passage times: proof of Theorem 2
We build upon ideas from [mcdiarmid1995minimal], while in contrast, we do not assume uniform boundedness of the increments. In the following, we write $C_{1},C_{2},C_{3},\dots$ as large constants and $c_{3},c_{4},\dots$ as small constants that may depend on the law of $\bm{\xi}$} and the branching rate. [^8] Not to be confused with the fixed constants $c_{1},c_{2}$ defined in Section 1.1.1. Here, we do not assume that $\bm{\xi}$} is spherically symmetric. We start with a few preparatory lemmas.
Lemma 11.
Assume (A6). For any $C_{1}>0$ and $c_{3}>0$, there exists $C_{2}>0$ such that the following holds for $k$ large enough: let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$}, $C>C_{2}$, and $(F_{Ck},\mathbf{J}_{Ck})=\sum_{j=1}^{Ck}(X_{j},\mathbf{Y}_{j})$. Then uniformly for $k$ large enough and ${\mathbf{z}}\in\mathbb{R}^{d}$ such that $\left\lVert{\mathbf{z}}\right\rVert\leqslant C_{1}k$,
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})\gg e^{-c_{3}k}.$
The proof of Lemma 11 is deferred to the appendix.
Lemma 12.
There exist some $C_{3},c_{4}>0$ such that for $k\geqslant 1$ large enough,
$\mathbb{P}(\#\{v\in V_{C_{3}k}:\eta^{(1)}_{v,C_{3}k}(C_{3}k)\geqslant k\}<e^{c_{4}k}\mid S)\ll e^{-c_{4}k}.$
Proof.
We first claim that for any $c_{5}>0$, there exist $c_{6},c_{7}>0$ such that uniformly for $n$ large enough,
$\mathbb{P}(\#\{v\in V_{n}:\left\lVert\bm{\eta}_{v,n}(n)\right\rVert\leqslant c_{5}n\}<e^{c_{6}n}\mid S)\ll e^{-c_{7}n}.$ (27)
Indeed, given $c_{5}$, let $c_{8}=c_{5}/(6d\bar{c}_{1})$ where here $\bar{c}_{1}$ stands for the maximum linear speed of the BRW across all $d$ directions, so that by Theorem 3.2 of [gantert2018large], [^9] Theorem 3.2 of [gantert2018large] is stated in the Schröder case, where $p_{0}+p_{1}>0$. However, the proof for the upper bound of the large deviation probability does not rely on this assumption; see Remark 3.3(b) therein.
$\mathbb{P}\left(\exists v\in V_{c_{8}n},\penalty 10000\ \left\lVert\bm{\eta}_{v,c_{8}n}(c_{8}n)\right\rVert>\frac{c_{5}n}{2}\right)\ll e^{-c_{7}n}.$ (28)
Conditioning upon the survival event $S$, with exponentially small probability we have that $V_{c_{8}n}\leqslant e^{c_{9}n}$. Evidently, each $v\in V_{c_{8}n}$ will lead independently with probability >1/2>1/2 to a particle $w\in V_{n}$ where $\left\lVert\bm{\eta}_{w,n}(n)-\bm{\eta}_{w,n}(c_{8}n)\right\rVert\leqslant c_{5}n/2$. This along with standard binomial estimates proves (27).
We return to the proof. We apply first (27) with $n=k$ and $c_{5}=1$. Outside a set of exponentially small probability, we may consider a subcollection of $e^{c_{6}k}$ particles in $\{v\in V_{k}:\left\lVert\bm{\eta}_{v,k}(k)\right\rVert\leqslant k\}$, which we label by $v_{j},\penalty 10000\ 1\leqslant j\leqslant e^{c_{6}k}$. We next evolve the particles $v_{j}$ independently and check if they reach $k$ in the $x$-direction. It suffices to consider the presence of the particles evolved from $v_{j},\penalty 10000\ 1\leqslant j\leqslant e^{c_{6}k}$ in a ball of radius one and at most $k+k+1\leqslant 3k$ apart (from the locations of $v_{j}$). By Lemma 11 applied with $C_{1}=3$ and $c_{3}=c_{6}/2$, for each $j$ we have for some $C_{2}>0$, with $w_{j}$ denoting the location of (a fixed descendant of) $v_{j}$ after time $C_{2}k$, that
$\mathbb{P}(w_{j}\in[k,\infty)\times\mathbb{R}^{d-1})\geqslant\mathbb{P}(w_{j}\in B_{(k+1,\mathbf{0})})\gg e^{-c_{6}k/2}.$ (29)
Note that such a fixed path exists for at least half of $v_{j}$’s with overwhelming probability. In conclusion, we evolved at least $e^{c_{6}k}/2$ many particles independently at time $k$ for time $C_{2}k$, each resulting in a probability of $\gg e^{-c_{6}k/2}$ to land in $k,\infty)\times\mathbb{R}^{d-1}$. A standard binomial estimate yields that with overwhelming probability, there exist $e^{c_{6}k/4}$ particles at time $(C_{2}+1)k$ present in $k,\infty)\times\mathbb{R}^{d-1}$ (conditional on survival). This completes the proof. ∎
Lemma 13.
There exist constants $c_{10},C_{4}>0$ such that for $x$ large enough,
$\mathbb{P}(\tau_{x}\geqslant C_{4}x\mid S)\ll e^{-c_{10}x}.$
Proof.
The proof is almost the same as Lemma 12, by replacing $k$ by $x$ and using the second inequality of (29) only. ∎
In the proof of Theorem 2, we wish to use Lemma 13 to bound probabilities of the form $\mathbb{P}(\tau_{x}\geqslant\alpha,\penalty 10000\ \tau_{y}\leqslant\beta)$, where $\beta<\alpha$. This is intuitively possible since we may run the particle $v_{y}$ that reaches $B_{y}$ and bound the probability that it never reaches $B_{x}$ in a time longer than $\alpha-\beta$, where we have used the strong Markov property of the random walk. There is, however, a subtle issue for the survival of the branching process initiated by the particle $v_{y}$, as this event is almost independent of the survival event $S$ we are conditioning on (in the proof of Theorem 2). We will bypass this difficulty by excluding this termination event a priori, or by first having an exponential number $e^{cn}$ of particles in $B_{y}$ instead of a single particle $v_{y}$.
Recall our notations beginning Section 2 and denote by $q=1-p=1-\mathbb{P}(S)$.
Proof of Theorem 2.
Pick $C_{3}$ and $c_{4}$ as in Lemma 12 and let $\delta\in(e^{-c_{13}/(2C_{3})},1)$ for some $c_{13}>0$ to be determined. For $x,y>0$, let
$s=s(x,y):=\inf\{t\geqslant 0:\mathbb{P}(\tau_{x}\leqslant t\mid S)\geqslant\delta^{y}\}.$
It follows that $\mathbb{P}(\tau_{x}<s-1\mid S)\leqslant\delta^{y}$ and
$\mathbb{P}(\tau_{x}>s+1)\leqslant 1-p\delta^{y}\leqslant\exp(-p\delta^{y}).$ (30)
Let $k=y/C_{3}$ and set $C_{5}\gg 1+C_{3}+C_{6}C_{7}$, all to be determined. Denote by $\tau^{(c_{11})}_{x+k}$ the waiting time until $e^{c_{11}k}$ many particles reach the ball $B_{x+k}$. Now note that
$\begin{split}\mathbb{P}(\tau_{x}>s-1)&\leqslant\mathbb{P}(s-1<\tau_{x}<s+C_{5}k)+\mathbb{P}\left(\tau^{(c_{11})}_{x+k}>s+C_{3}k+1+C_{6}C_{7}k\right)\\ &\hskip 71.13188pt+\mathbb{P}\left(\tau_{x}\geqslant s+C_{5}k,\penalty 10000\ \tau^{(c_{11})}_{x+k}\leqslant s+C_{3}k+1+C_{6}C_{7}k\right).\end{split}$ (31)
We first decompose the second event $\tau^{(c_{11})}_{x+k}>s+C_{3}k+1+C_{6}C_{7}k$ of (31) depending on the number of particles at time $C_{3}k$ that lie beyond $k$ in the $x$-coordinate. On one hand, by Lemma 12,
$\mathbb{P}(\#\{v\in V_{C_{3}k}:\eta^{(1)}_{v,C_{3}k}(C_{3}k)\geqslant k\}<e^{c_{4}k})\leqslant q+o(e^{-c_{4}k}).$
On the other hand, on the event that $\{v\in V_{C_{3}k}:\eta^{(1)}_{v,C_{3}k}(C_{3}k)\geqslant k\}\geqslant e^{c_{4}k}$, we may label $e^{c_{4}k}$ many particles by $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$. Our goal is to show that the probability that fewer than $e^{c_{11}k}$ of $\{v_{i}\}_{1\leqslant i\leqslant e^{c_{4}k}}$ reach $B_{x}$ in the period of time $C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ is small. Note that the trajectories for each $i$ in the period $C_{3}k,C_{3}k+s+1+C_{6}C_{7}k]$ are independent and equivalent in law to a BRW trajectory initiated at $\bm{\eta}_{v,C_{3}k}(C_{3}k)$. A technical difficulty arises, however, as the events of reaching $B_{x}$ are not monotone in $x$ in dimensions $d\geqslant 2$. It would be nice if we had stochastic monotonicity for $\tau_{x}$ in $x$ (as in the one-dimensional case), but this appears out of our reach. Instead, we use the following argument to bypass this technical difficulty.
Recall that each $\eta^{(1)}_{v_{i},C_{3}k}(C_{3}k)\geqslant k$, and the notation $\mathbf{x}=(x,0\dots,0)\in\mathbb{R}^{d}$. By a standard large deviation estimate on the last d1d-1 coordinates of $\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)$ and using the triangle inequality, we may assume that $\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k$ for all $1\leqslant i\leqslant e^{c_{4}k}$ by losing an event of probability $O(e^{-c_{12}k})$ with $c_{4}$ chosen small enough. Similarly, we may assume $\eta^{(1)}_{v_{i},C_{3}k}(C_{3}k)\leqslant C_{8}k$ with an overwhelming probability $1-O(e^{-c_{12}k})$ (since there are $e^{c_{4}k}$ many particles in total and we can apply a union bound). As a consequence,
$\mathbb{P}\left(\forall 1\leqslant i\leqslant e^{c_{4}k},\penalty 10000\ \max\{0,x-C_{8}k\}\leqslant\left\lVert\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert\leqslant x+C_{6}k\right)\geqslant 1-O(e^{-c_{12}k}).$
For each $1\leqslant i\leqslant e^{c_{4}k}$, consider the set of $\mathbf{x}_{i}$’s such that $\left\lVert\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)\right\rVert=x$. Selecting $C_{8}>C_{6}$ and on a set of overwhelming probability of at least $1-O(e^{-c_{12}k})$, we can select $\mathbf{x}_{i}$ such that $\left\lVert\mathbf{x}-\mathbf{x}_{i}\right\rVert\leqslant C_{8}k$. For ${\mathbf{z}}\in\mathbb{R}^{d}$, we denote by $\tau_{{\mathbf{z}}}$ the FPT to the ball of radius one centered at $\mathbf{z}$}. By Lemma 13 applied with $C_{4}=C_{7}/2$ and picking $C_{7}$ large (and without loss of generality, a $p/2$ proportion of the $e^{c_{4}k}$ many particles do not terminate, by losing an exponentially small probability),
$\hskip 14.22636pt\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k\}\geqslant e^{c_{11}k}\right)$
$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)+O(e^{-c_{12}k})$
$\hskip 28.45274pt+\mathbb{P}\left(\exists 1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1+C_{6}C_{7}k,\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}<s+1\right)$
$\leqslant\mathbb{P}\left(\#\{1\leqslant i\leqslant\frac{p}{2}e^{c_{4}k},\penalty 10000\ \tau_{\mathbf{x}_{i}-\bm{\eta}_{v_{i},C_{3}k}(C_{3}k)}\geqslant s+1\}\geqslant e^{c_{11}k}\right)$
$\hskip 113.81102pt+\sum_{1\leqslant i\leqslant e^{c_{4}k}}\mathbb{P}\left(\tau^{(i)}_{C_{6}k+1}>C_{6}C_{7}k\right)+O(e^{-c_{12}k})$
$\leqslant\mathbb{P}\left(\mathrm{Bin}(\frac{p}{2}e^{c_{4}k},\mathbb{P}(\tau_{x}>s+1))\geqslant e^{c_{11}k}\right)+e^{c_{4}k}e^{-c_{10}k}+O(e^{-c_{12}k}).$
Here and later, we use $\mathrm{Bin}(n,p)$ to denote a generic binomial random variable with $n$ trials and success probability $p$. By choosing $c_{4}$ and then $c_{11}$ small enough and using (30), we obtain that the above satisfies
$\mathbb{P}\left(\mathrm{Bin}\left(\frac{p}{2}e^{c_{4}k},\mathbb{P}(\tau_{x}>s+1)\right)\geqslant e^{c_{11}k}\right)+e^{(c_{4}-c_{10})k}+O(e^{-c_{12}k})$
$\hskip 199.16928pt\leqslant\exp(-p\delta^{y}e^{c_{13}k})+O(e^{-c_{14}k}).$
Inserting back to (31) and using Lemma 13 on the third term on the right-hand side of (31) (with our choice of $C_{5}\gg 1+C_{3}+C_{6}C_{7}$), we have
$\mathbb{P}(\tau_{x}>s-1)\leqslant q+\exp(-p\delta^{y}e^{c_{13}k})+O(e^{-c_{14}k})+\mathbb{P}(s-1<\tau_{x}\leqslant s+C_{5}k).$ (32)
Since $\delta\in(e^{-c_{13}/(2C_{3})},1)$ and $k=y/C_{3}$, for $y$ large enough we have $\exp(-p\delta^{y}e^{c_{13}k})\leqslant\delta^{y}$. Thus, we conclude from (32) that
$\mathbb{P}(\tau_{x}>s+C_{5}k)\leqslant q+O(e^{-c_{15}y}).$ (33)
For $q>0$, using the fact that $\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$ (this is a consequence of Theorem 13.3 of [athreya2004branching]), we obtain
$\begin{split}\mathbb{P}(\tau_{x}>s+C_{5}k)&\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q\,\mathbb{P}(\tau_{x}>y,\penalty 10000\ \#V_{y}=0\mid S^{c})\\ &\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q(1-o(e^{-c_{16}y}))-C_{9}\mathbb{P}(\tau_{x}\leqslant y,\penalty 10000\ S^{c}).\end{split}$ (34)
Again by using $\mathbb{P}(\#V_{n}>0\mid S^{c})=o(e^{-c_{16}n})$, we have for some $c_{17}$ to be determined later,
$\mathbb{P}(\tau_{x}\leqslant y,\penalty 10000\ S^{c})\leqslant o(e^{-c_{17}c_{16}y})+\mathbb{P}(\tau_{x}\leqslant c_{17}y).$
Recall that $M_{n}$ denotes the maximum in the first direction of the BRW at time step $n$. Using Lemma 5.2 of [gantert2018large], it holds that for $\widetilde{M}_{j}$ denoting the maxima of $V_{j}$ many independent random walks,
$\mathbb{P}(\tau_{x}\leqslant c_{17}y)\leqslant\sum_{j=1}^{c_{17}y}\mathbb{P}(M_{j}\geqslant x)\leqslant\sum_{j=1}^{c_{17}y}\mathbb{P}(\widetilde{M}_{j}\geqslant x).$
It is straightforward to show using a Chebyshev argument that uniformly for $y\leqslant x$, with $c_{17}$ small enough, the right-hand side is bounded by $e^{-c_{18}y}$. Inserting into (34) yields
$\mathbb{P}(\tau_{x}>s+C_{5}k)\geqslant p\,\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)+q-C_{10}e^{-c_{19}y}.$
Using (33), it then holds that
$\mathbb{P}(\tau_{x}>s+C_{5}k\mid S)\ll e^{-c_{19}y}.$
We conclude that for $y\leqslant x$ large enough,
$\mathbb{P}(s-1\leqslant\tau_{x}\leqslant s+C_{5}k\mid S)\geqslant 1-\delta^{y}-O(e^{-c_{19}y})\geqslant 1-O(e^{-c_{20}y}).$
Therefore, for some $y_{0}>0$ large enough, $\mathbb{P}(s-1\leqslant\tau_{x}\leqslant s+C_{5}y_{0}/C_{3}\mid S)>4/5$, so that $s-1\leqslant t^{(1/2)}_{x}\leqslant s+C_{5}y_{0}/C_{3}$. For $y>2C_{5}y_{0}/C_{3}$, we then have
$\mathbb{P}\left(|\tau_{x}-t^{(1/2)}_{x}|>y\mid S\right)$ $\leqslant 1-\mathbb{P}\left(s-1\leqslant\tau_{x}\leqslant s+\frac{y}{2}+\frac{C_{5}y_{0}}{C_{3}}\mid S\right)\leqslant O(e^{-c_{20}C_{3}y/(2C_{5})}).$
For the case $y\leqslant 2C_{5}y_{0}/C_{3}$, we may increase the constant $C$ in (5) such that for $y$ in this range, $Ce^{-cy}\geqslant 1$ holds, and hence (5) trivially holds. This establishes (5). ∎
We remark that the above proof works for a general $r\in(0,1)$ (where $r$ is the quantile in $t_{x}^{(r)}$), in addition to the case $r=1/2$.
2.4 Proof of Theorem 3
First, let us recall our notations. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel–Legendre transform, which coincides with the large deviation rate function for the jump distribution $\bm{\xi}$}. The unique positive solution to $I(\widehat{c}_{1},\mathbf{0})=\log\rho$ is given by $\widehat{c}_{1}$. Our goal is to prove that $\tau_{x}=\widehat{A}(x)+o_{\mathbb{P}}(\log x)$, where
$\widehat{A}(x):=\frac{x}{\widehat{c}_{1}}+\left(\frac{d+2}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\right)\log x.$
Figure 3: Visualization of how random walks in cones are related to estimating first passage times. The first passage time $\tau_{x}$ is roughly the time when the growing range (illustrated with the shaded ellipses) of the BRW becomes tangent with the target ball $B_{x}$. In our proof, we will construct moving cones (that are circular centered in the direction $\mathbf{c}_{2}$ in the sense of (35); the boundaries are indicated by the blue lines) and show that with high probability, the range of the BRW always lies within the moving cones.
2.4.1 Random walks in cones
For estimating $\tau_{x}$, we need a few preliminary results on random walks and Brownian motion in cones. The intuitive reason why random walks in cones come into play is explained in Figure 3. In this section, we try to be self-contained and summarize only the necessary results. We refer to [banuelos1997brownian, burkholder1977exit, deblassie1987exit, denisov2015random] for more background on this topic.
We first fix some notations. Let $\{\mathbf{U}_{i}\}_{i\geqslant 1}$ be i.i.d. copies of $\mathbf{U}=(U_{1},\dots,U_{d})$ in $\mathbb{R}^{d}$ and $\mathbf{S}_{n}=\mathbf{U}_{1}+\dots+\mathbf{U}_{n}$ denote its partial sum. Consider an open connected subset $\mathfrak{S}\subseteq\mathbb{R}^{d}$ and let $\mathbb{K}$ be the cone generated by the rays emanating from $\mathbf{0}\in\mathbb{R}^{d}$ that go through $\mathfrak{S}$. For ${\mathbf{z}}\in\mathbb{K}$, let
$\tau^{\mathbb{K}}_{\mathbf{z}}=\min\{n\geqslant 1:{\mathbf{z}}+\mathbf{S}_{n}\not\in\mathbb{K}\}$
be the first passage time of the random walk $\{\mathbf{S}_{n}\}_{n\geqslant 1}$ to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. Similarly, let $\tau_{\mathbf{z}}^{\mathbb{K},\mathrm{BM}}$ be the first passage time of the standard $d$-dimensional Brownian motion to $\mathbb{K}^{\mathrm{c}}$, starting from ${\mathbf{z}}\in\mathbb{K}$. We will frequently consider cones that are circular (in the sense of [deblassie1987exit]), that is, of the form
$\mathbb{K}_{\alpha}(\mathbf{v}):=\{{\mathbf{z}}\in\mathbb{R}^{d}:0\leqslant\theta<\alpha\},$ (35)
where $\alpha\in(0,\pi),\mathbf{0}\neq\mathbf{v}\in\mathbb{R}^{d}$, and $\theta=\theta({\mathbf{z}},\mathbf{v})$ is the angle between nonzero vectors $\mathbf{z}$} and $\mathbf{v}$ in $\mathbb{R}^{d}$. In the case $\mathbf{v}=(1,\mathbf{0})$, we may write $\mathbb{K}_{\alpha}=\mathbb{K}_{\alpha}((1,\mathbf{0}))$.
The following result combines several results of [denisov2015random]. Recall from assumption (A5) that the law of $\bm{\xi}$} is non-lattice.
Lemma 14.
Fix a circular cone $\mathbb{K}$, i.e., $\mathbb{K}=\mathbb{K}_{\alpha}$ for some $\alpha\in(0,\pi)$. Suppose that $\mathbf{U}$ is centered, non-lattice, and $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$. Let $\mathbf{x},\mathbf{y}\in\mathbb{K}$. There exist positive functions $V,\widetilde{V}$ satisfying $\max(V(\mathbf{x}),\widetilde{V}(\mathbf{x}))\leqslant C(1+\left\lVert\mathbf{x}\right\rVert)^{p}$ and
$\mathbb{P}(\mathbf{x}+\mathbf{S}_{n}\in B_{\mathbf{y}},\tau^{\mathbb{K}}_{\mathbf{x}}>n)\ll\frac{V(\mathbf{x})\widetilde{V}(\mathbf{y})}{n^{p+d/2}},$ (36)
where $p=p(\mathbb{K})$ is such that $\mathbb{P}(\tau_{\mathbf{x}}^{\mathbb{K},\mathrm{BM}}>t)\asymp t^{-p/2}$. Moreover, for fixed $\mathbf{y}\in\mathbb{K}$,
$\mathbb{P}(\mathbf{x}+\mathbf{S}_{n}\in B_{\mathbf{y}},\tau^{\mathbb{K}}_{\mathbf{x}}>n)\asymp\frac{V(\mathbf{x})}{n^{p+d/2}},$ (37)
where the asymptotic constant in $\asymp$ may depend on $y$.
Proof.
The asymptotic (37) is a direct consequence of (1), (10) of Theorem 6 (which holds also for a fixed $\mathbf{y}\in\mathbb{K}$), and Lemma 13(b) of [denisov2015random]. On the other hand, (36) follows from the derivation of Lemmas 27 and 28 therein, along with their Theorem 1. The results in [denisov2015random] are stated in the lattice case but non-lattice walks can be proved in the same way. ∎
Remark 2.
As commented in [denisov2015random], if $\mathbf{U}$ does not satisfy $\mathbb{E}[U_{i}U_{j}]={\mathbbm{1}}_{\{i=j\}}$ but has a positive definite covariance matrix, then the same results hold with $\mathbb{K}$ replaced by $M\mathbb{K}=\{M\mathbf{x}:\mathbf{x}\in\mathbb{K}\}$, where $\mathbf{V}=M\mathbf{U}$ satisfies $\mathbb{E}[V_{i}V_{j}]={\mathbbm{1}}_{\{i=j\}}$.
The exponent $p(\mathbb{K})$ is in general not explicit and is related to the smallest eigenvalue of the Laplace-Beltrami operator; see [denisov2015random]. However, it suffices for our purpose to understand the behavior of $p(\mathbb{K})$ when $\mathbb{K}$ is a circular cone that is close to a half-space. The following result shows that $p\approx 1$ if the circular cone $\mathbb{K}$ is close enough to a half-space.
Lemma 15 ((3.11) of [burkholder1977exit]).
Let $\alpha\in(0,\pi)$ and ${\mathbf{z}}\in\mathbb{K}_{\alpha}$ be arbitrary. There exists a decreasing and continuous function $p_{\alpha,d}$ such that $p_{\pi/2,d}=1$ and [^10] The formal definition is given in [burkholder1977exit] by $p_{\alpha,d}=\sup\{p:\alpha<\theta_{p,d}\}$, where $\theta_{p,d}$ is the smallest zero in $(0,\pi)$ of $h(\theta)=F(-p,p+d-2;(d-1)/2;(1-\cos\theta)/2)$. The continuity is not explicitly stated in [burkholder1977exit], but follows from the continuity of the hypergeometric function.
$\mathbb{E}[(\tau^{\mathbb{K}_{\alpha}}_{\mathbf{z}})^{q/2}]<\infty\quad\Longleftrightarrow\quad q<p_{\alpha,d}.$
The following ballot-type lemma provides bounds when the cone is shifted by a function of time. We do not attempt the sharpest bound here, but state the minimal result needed for our purpose.
Lemma 16.
Assume the same setting on the law of $\mathbf{U}$ as in Lemma 14. Let $\mathbf{v}\in\mathbb{R}^{d},\penalty 10000\ \mathbf{v}\neq\mathbf{0}$, and $\alpha\in(0,\pi)$. Let $p=p(\mathbb{K}_{\alpha}(-\mathbf{v}))$ be the exponent defined in Lemma 14, $p^{\prime}<p$, and $C>0$. Then uniformly in ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ such that ${\mathbf{z}}-\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{v})$ and $n\in\mathbb{N}$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{n}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant n,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}},\end{split}$ (38)
where the asymptotic constant in $\ll$ depends on $d,p^{\prime}$, and the law of $\bm{\xi}$}. Moreover, uniformly for $n\in\mathbb{N}$, $1\leqslant\ell\leqslant n$, and ${\mathbf{z}},\mathbf{y}\in\mathbb{R}^{d}$ with ${\mathbf{z}}-\mathbf{y}\in C\log(\min(\ell,n-\ell))_{+}\mathbf{v}+\mathbb{K}_{\alpha}(-\mathbf{v})$,
$\begin{split}&\mathbb{P}(\mathbf{S}_{\ell}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant\ell,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))\\ &\hskip 170.71652pt\ll\frac{(1+\left\lVert{\mathbf{z}}\right\rVert)^{p}(1+\left\lVert\mathbf{y}\right\rVert)^{p}}{n^{p^{\prime}+d/2}},\end{split}$ (39)
where the asymptotic constant in $\ll$ depends on $d,p^{\prime}$, and the law of $\bm{\xi}$}.
Proof.
Let $\tilde{\mathbf{y}}=-\mathbf{y}-(C\log n)\mathbf{v}$ and $\tilde{{\mathbf{z}}}=-{\mathbf{z}}-(C\log n)\mathbf{v}$. We have by using (36) of Lemma 14,
$\hskip 14.22636pt\mathbb{P}(\mathbf{S}_{n}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant n,\,\mathbf{S}_{k}-\mathbf{y}-C\log(\min(k,n-k))_{+}\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))$
$\leqslant\mathbb{P}(\mathbf{S}_{n}\in B_{\mathbf{z}};\,\forall 1\leqslant k\leqslant n,\,\mathbf{S}_{k}-\mathbf{y}-(C\log n)\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{v}))$
$=\mathbb{P}(\tilde{\mathbf{y}}+\mathbf{S}_{n}\in B_{\tilde{{\mathbf{z}}}};\,\forall 1\leqslant k\leqslant n,\,\tilde{\mathbf{y}}+\mathbf{S}_{k}\in\mathbb{K}_{\alpha}(-\mathbf{v}))$
$\ll\frac{(1+\left\lVert\tilde{\mathbf{y}}\right\rVert)^{p}(1+\left\lVert\tilde{{\mathbf{z}}}\right\rVert)^{p}}{n^{p+d/2}}$
$\ll\frac{(1+\left\lVert{\mathbf{y}}\right\rVert)^{p}(1+\left\lVert{{\mathbf{z}}}\right\rVert)^{p}}{n^{p^{\prime}+d/2}}.$
This completes the proof of (LABEL:eq:ballot1). The proof of (LABEL:eq:ballot2) is similar and is omitted here for brevity. ∎
2.4.2 A large deviation computation
In the following, we work with a sequence $\{{\bm{\xi}}_{j}=(X_{j},\mathbf{Y}_{j})\}_{j\in\mathbb{N}}$ of i.i.d. centered random variables in $\mathbb{R}^{d}$ that are not necessarily spherically symmetric but have densities belonging to $L^{r}(\mathbb{R}^{d})$ for some $r>1$ and a positive definite covariance matrix $V$. Let $\Lambda(\bm{\lambda})=\log\phi_{\bm{\xi}}(\bm{\lambda})=\log\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ and $I(\mathbf{x})$ be its Fenchel–Legendre transform. [^11] Not to be confused with the functions $\Lambda$ and II for the first coordinate which is defined on $\mathbb{R}$, since we may assume $d\geqslant 2$. Let also $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})$. By definition,
$\Lambda(\mathbf{c}_{2})=\mathbf{c}_{2}\cdot\left(\frac{x}{n},\mathbf{0}\right)-I\left(\frac{x}{n},\mathbf{0}\right).$ (40)
We have the following counterpart of Proposition 10. The proof is similar and deferred to the appendix.
Proposition 17.
Fix a large constant $C>0$ and a dimension $d\geqslant 1$. Suppose that the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$} satisfies conditions (A2), (A5), and (A6) from Section 1.1. Let $(X_{j},\mathbf{Y}_{j})$ be a sequence of i.i.d. random variables with the same distribution as $\bm{\xi}$} and $(F_{n},\mathbf{J}_{n})=\sum_{j=1}^{n}(X_{j},\mathbf{Y}_{j})$. Fix $\overline{c}>\underline{c}>0$.
  1. (i)
    Uniformly for any $c=c(n)\in[-C\log n,C\log n]$ and any $x\in[\underline{c}\,n,\overline{c}\,n]$,
    $\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\mid F_{n}\geqslant x+c)\ll e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{0})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.$ (41)
  2. (ii)
    Fix $a(n)\to\infty$ and $a(n)=o(\log n)$. Uniformly for $\mathbf{y}$ with $\left\lVert\mathbf{y}\right\rVert\leqslant C\sqrt{n}$ and $c\in[-C\log n,C\log n]$, and any $x\in[\underline{c}\,n,\overline{c}\,n]$,
    $\begin{split}&\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1)\mid F_{n}\in[x+c,x+c+a(n)])\\ &\hskip 56.9055pt\gg e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{y})}n^{-(d-1)/2}e^{nI(x/n)+cI^{\prime}(x/n)}.\end{split}$ (42)
2.4.3 Upper bound of Theorem 3
Let $\varepsilon>0$ be arbitrary and
$n=\widehat{A}_{1}(x)=\frac{x}{\widehat{c}_{1}}+\left(\frac{d+2+2\varepsilon}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\right)\,\log x.$ (43)
Our goal is to show that the event $\{\tau_{x}>\widehat{A}_{1}(x)\}$ has a small probability. Define $\mathbf{c}_{2}:=\nabla I(x/n,\mathbf{0})$. By definition,
$\Lambda(\mathbf{c}_{2})=\mathbf{c}_{2}\cdot\left(\frac{x}{n},\mathbf{0}\right)-I\left(\frac{x}{n},\mathbf{0}\right).$ (44)
We set up a barrier event. Pick $\alpha\in(0,\pi/2)$ such that $p_{\alpha,d}=1+\varepsilon$; cf., Lemma 15. Recall (35). Let
$\widehat{P}_{n}:=\left\{v\in V_{n}:\penalty 10000\ \forall 0\leqslant k\leqslant n,\penalty 10000\ \bm{\eta}_{v,n}(k)-\frac{k\mathbf{x}}{n}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right\}$
and let
$\widehat{\Delta}_{n,x}:=\sum_{v\in\widehat{P}_{n}}{\mathbbm{1}}_{\widehat{H}_{v,n}(x)}:=\sum_{v\in\widehat{P}_{n}}{\mathbbm{1}}_{\{\bm{\eta}_{v,n}(n)\in B_{x}\}}$ (45)
be the number of particles lying in the target $B_{x}$ at time $n=\widehat{A}_{1}(x)$ that is controlled by the barrier given by $\widehat{P}_{n}$. The quantity $\widehat{\Delta}_{n,x}$ serves as a lower bound for the number of particles lying in the target $B_{x}$ at time $n$. As a consequence, it suffices to show that
$\mathbb{P}(\widehat{\Delta}_{n,x}\geqslant 1)>1/C$ (46)
for some $C>0$ independent of $x$, which in turn gives the upper bound for $\tau_{x}$ by Theorem 2. We compute first and second moments for $\widehat{\Delta}_{n,x}$. Applying the many-to-one formula, the change of measure (72), and using a similar change of measure argument as in the proof of Proposition 17, we have
$\mathbb{E}[\widehat{\Delta}_{n,x}]$ $=\rho^{n}\mathbb{P}\left(\forall 0\leqslant k\leqslant n,\penalty 10000\ \mathbf{S}_{k}-\frac{k\mathbf{x}}{n}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\mathbf{S}_{n}\in B_{x}\right)$
$\gg\rho^{n}e^{-nI(x/n,\mathbf{0})}\mathbb{Q}\left(\forall 0\leqslant k\leqslant n,\mathbf{S}_{k}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\left\lVert\mathbf{S}_{n}\right\rVert\leqslant 1\right).$
By (37) of Lemma 14,
$\mathbb{Q}\left(\forall 0\leqslant k\leqslant n,\mathbf{S}_{k}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2});\,\left\lVert\mathbf{S}_{n}\right\rVert\leqslant 1\right)\asymp n^{-p_{\alpha,d}-d/2}\asymp x^{-p_{\alpha,d}-d/2}.$
Here, we may without loss of generality replace $\mathbf{c}_{2}$ by $M\mathbf{c}_{2}$ where $\mathbb{E}_{\mathbb{Q}}[(M{\bm{\xi}})_{i}(M{\bm{\xi}})_{j}]={\mathbbm{1}}_{\{i=j\}}$. This does not affect the analysis since an invertible linear transformation of a half-space containing $\mathbf{0}$ is a half-space containing $\mathbf{0}$ and we are in a regime where $p_{\alpha,d}\approx 1$; see Remark 2. Using (43), we have the approximation
${nI\Big(\frac{x}{n},\mathbf{0}\Big)}=\frac{\log\rho}{\widehat{c}_{1}}\,x+\left(\frac{(d+2+2\varepsilon)\log\rho}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}-(d+2+2\varepsilon)\right)\log x+O(1).$
These altogether lead to
$\mathbb{E}[\widehat{\Delta}_{n,x}]$ $\gg\rho^{n}\rho^{-x/\widehat{c}_{1}}\rho^{-{(d+2+2\varepsilon)\log x}/({2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})})}x^{(d+2+2\varepsilon)/2}x^{-p_{\alpha,d}-d/2}=1,$ (47)
where we used $p_{\alpha,d}=1+\varepsilon$.
Next, we compute the second moment of $\widehat{\Delta}_{n,x}$, following the same type of arguments that lead to (19). Recall the definition of $\widehat{H}_{v,n}(x)$ from (45). By conditioning on the closest integer point to $\bm{\eta}_{v,n}(n-s)\in(n-s)\mathbf{x}/n+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$ and applying (36) of Lemma 14, $\text{ for }v\sim_{s}w$ (recall that this means that the distance of $v$ and $w$ in the genealogical tree is equal to $2s$),
$\hskip 14.22636pt\mathbb{P}\left(\widehat{H}_{v,n}(x)\cap\widehat{H}_{w,n}(x)\right)$
$\ll\sum_{\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\cap\mathbb{Z}^{d}}\left(e^{-(n-s)I({x}/{n},\mathbf{0})-\mathbf{c}_{2}\cdot\mathbf{y}}(\max(1,n-s))^{-p_{\alpha,d}-d/2}(1+\left\lVert\mathbf{y}\right\rVert)^{2p_{\alpha,d}}\right)$
$\hskip 56.9055pt\times\left(e^{-sI({x}/{n},\mathbf{0})-\mathbf{c}_{2}\cdot(-\mathbf{y})}s^{-p_{\alpha,d}-d/2}(1+\left\lVert\mathbf{y}\right\rVert)^{2p_{\alpha,d}}\right)^{2}$
$\ll e^{-(n+s)I({x}/{n},\mathbf{0})}(\max(1,n-s))^{-p_{\alpha,d}-d/2}s^{-2p_{\alpha,d}-d}\sum_{\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\cap\mathbb{Z}^{d}}e^{\mathbf{c}_{2}\cdot\mathbf{y}}(1+\left\lVert\mathbf{y}\right\rVert)^{6p_{\alpha,d}}.$
By the definition (35),
$\sum_{\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\cap\mathbb{Z}^{d}}e^{\mathbf{c}_{2}\cdot\mathbf{y}}(1+\left\lVert\mathbf{y}\right\rVert)^{6p_{\alpha,d}}$ $=\sum_{j=0}^{\infty}\sum_{\begin{subarray}{c}\mathbf{y}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\cap\mathbb{Z}^{d}\\ \mathbf{y}\cdot(-\mathbf{c}_{2})\in[j,j+1]\end{subarray}}e^{\mathbf{c}_{2}\cdot\mathbf{y}}(1+\left\lVert\mathbf{y}\right\rVert)^{6p_{\alpha,d}}$
$\ll\sum_{j=0}^{\infty}e^{-j}(1+j)^{12dp_{\alpha,d}}\ll 1.$
As a consequence and using $p_{\alpha,d}=1+\varepsilon$, $\text{ for }v\sim_{s}w$,
$\mathbb{P}\left(\widehat{H}_{v,n}(x)\cap\widehat{H}_{w,n}(x)\right)$ $\ll e^{-(n+s)I({x}/{n},\mathbf{0})}(\max(1,n-s))^{-p_{\alpha,d}-d/2}s^{-2p_{\alpha,d}-d}$
$\ll\rho^{-(n+s)}x^{(p_{\alpha,d}+d/2)(n+s)/n}(\max(1,n-s))^{-p_{\alpha,d}-d/2}s^{-2p_{\alpha,d}-d}.$
Therefore, by assumption (A1),
$\mathbb{E}[\widehat{\Delta}_{n,x}^{2}]$ $\ll\mathbb{E}[\widehat{\Delta}_{n,x}]+\sum_{s=1}^{n}\rho^{n+s}\mathbb{P}\left(\widehat{H}_{v,n}(x)\cap\widehat{H}_{w,n}(x)\text{ for }v\sim_{s}w\right)$
$\ll 1+\sum_{s=1}^{n}x^{(p_{\alpha,d}+d/2)(n+s)/n}(\max(1,n-s))^{-p_{\alpha,d}-d/2}s^{-2p_{\alpha,d}-d}\ll 1,$
where in the last step we argue similarly as in the proof of Lemma 2.7 of [bramson2016convergence]. Combined with (47) and the Paley-Zygmund inequality leads to (46). This proves the desired upper bound.
2.4.4 Lower bound of Theorem 3
Fix $\varepsilon>0$. Define
$\widehat{A}_{2}(x):=\frac{x}{\widehat{c}_{1}}+\frac{(d+2-2\varepsilon)\log x}{2\,\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}.$
Our goal is to show that the event $\{\tau_{x}\leqslant\widehat{A}_{2}(x)\}$ has a small probability. We first consider the number $M_{x}$ of particles in $B_{x}$ at time $n=\widehat{A}_{2}(x)$. We find $\alpha\in(0,\pi)$ such that $p_{\alpha,d}=1-\varepsilon/2$. Recalling $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})\in\mathbb{R}_{+}\times\mathbb{R}^{d-1}$, we may assume that $\mathbf{x}/n\not\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$. We may decompose $M_{x}=\sum_{\ell=0}^{n}M_{x,\ell}$, where
$M_{x,n}:=\#\left\{v\in V_{n}:\bm{\eta}_{v,n}(n)\in B_{x};\,\forall 0\leqslant k\leqslant n,\,\bm{\eta}_{v,n}(k)\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right\},$
where $\bm{\psi}_{n}(k):=(C_{1}+C_{2}(\log(\min(k,n-k)))_{+})\,\mathbf{c}_{2}$ for some $C_{1},C_{2}>0$ to be determined, and for $0\leqslant\ell\leqslant n-1$,
$M_{x,\ell}:=\#\Bigg\{v\in V_{n}:\bm{\eta}_{v,n}(n)\in B_{x};\,$ $\forall 0\leqslant k\leqslant\ell,\,\bm{\eta}_{v,n}(k)\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),$
$\bm{\eta}_{v,n}(\ell+1)\not\in\frac{(\ell+1)\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\Bigg\}.$
Denote by $\{\mathbf{S}_{n}\}_{n\geqslant 1}$ the random walk with i.i.d. increments $\{{\bm{\xi}}_{i}\}_{i\geqslant 1}$. Applying the change of measure (72) and (LABEL:eq:ballot1) of Lemma 16, we have
$\mathbb{E}[M_{x,n}]$ $=\rho^{n}\mathbb{P}\left(\mathbf{S}_{n}\in B_{x};\,\forall 0\leqslant k\leqslant n,\,\mathbf{S}_{k}\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right)$
$=\rho^{n}e^{-nI(x/n,\mathbf{0})}\mathbb{Q}\left(\mathbf{S}_{n}\in B_{\mathbf{0}};\,\forall 0\leqslant k\leqslant n,\,\mathbf{S}_{k}\in\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\right)$
$\ll\rho^{n}e^{-nI(x/n,\mathbf{0})}n^{-p_{\alpha,d}+\varepsilon/4-d/2}(1+C_{1})^{p_{\alpha,d}}$
$\ll x^{-\varepsilon/4}.$
Next, we bound $\mathbb{E}[M_{x,\ell}]$ for $0\leqslant\ell\leqslant n-1$. We condition on the location of $\bm{\eta}_{v,n}(\ell)$, which may take value in ${\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$. This leads to
$\begin{split}&\hskip 14.22636pt\mathbb{E}[M_{x,\ell}]\\ &\leqslant\rho^{\ell}\sum_{\begin{subarray}{c}\mathbf{u}\in\mathbb{Z}^{d}\\ \mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\end{subarray}}\mathbb{P}\left(\forall 0\leqslant k\leqslant\ell,\,\mathbf{S}_{k}\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),\,\mathbf{S}_{\ell}\in B_{\mathbf{u}}\right)\\ &\hskip 128.0374pt\times\mathbb{P}\left({\bm{\xi}}\in-\mathbf{u}+\frac{(\ell+1)\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)+(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}\right).\end{split}$ (48)
We bound separately the above two probabilities for a fixed $\ell$. Applying the change of measure (72) and (LABEL:eq:ballot2) of Lemma 16,
$\hskip 14.22636pt\mathbb{P}\left(\forall 0\leqslant k\leqslant\ell,\,\mathbf{S}_{k}\in\frac{k\mathbf{x}}{n}+\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),\,\mathbf{S}_{\ell}\in B_{\mathbf{u}}\right)$
$=e^{-\ell I(\mathbf{u}/\ell)}\mathbb{Q}\left(\forall 0\leqslant k\leqslant\ell,\,\mathbf{S}_{k}\in\bm{\psi}_{n}(k)+\mathbb{K}_{\alpha}(-\mathbf{c}_{2}),\,\mathbf{S}_{\ell}\in B_{\mathbf{u}-\ell\mathbf{x}/n}\right).$
$\ll e^{-\ell I(\mathbf{u}/\ell)}n^{-p_{\alpha,d}+\varepsilon/4-d/2}(1+C_{1})^{p_{\alpha,d}}$
$\ll\rho^{-\ell}x^{-\varepsilon/4}\left\lVert\mathbf{u}-\frac{\ell\mathbf{x}}{n}\right\rVert e^{-\mathbf{c}_{2}\cdot(\mathbf{u}-\ell\mathbf{x}/n)}.$
On the other hand, since the moment generating function of $\bm{\xi}$} exists in a neighborhood of $\mathbf{c}_{2}$, we have for some $\delta>0$,
$\mathbb{P}({\bm{\xi}}\in B_{x})\ll e^{-(1+\delta)\mathbf{c}_{2}\cdot\mathbf{x}}.$
It follows that for $\mathbf{v}\in\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$,
$\hskip 14.22636pt\mathbb{P}\left({\bm{\xi}}\in-\mathbf{v}+\frac{\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)-\bm{\psi}_{n}(\ell)+(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}\right)$
$\ll\exp\left(-(1+\delta)\inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}+\frac{\mathbf{x}}{n}+\bm{\psi}_{n}(\ell+1)-\bm{\psi}_{n}(\ell)\right)\right)\right)$
$\leqslant\exp\left(-(1+\delta)\inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\right),$
where we have used that $(x/n,\mathbf{0})\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{c}$.
Inserting the above estimates back to (LABEL:eq:Mxl) with the change of variable $\mathbf{v}=\mathbf{u}-({\ell\mathbf{x}}/{n}+\bm{\psi}_{n}(\ell))$ leads to
$\mathbb{E}[M_{x,\ell}]$ $\ll\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\end{subarray}}x^{-\varepsilon/4}\left\lVert\mathbf{v}+\bm{\psi}_{n}(\ell)\right\rVert e^{-\mathbf{c}_{2}\cdot(\mathbf{v}+\bm{\psi}_{n}(\ell))}$
$\hskip 85.35826pt\times\exp\left(-(1+\delta)\inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\right)$
$\ll x^{-\varepsilon/4}\sum_{j=0}^{\infty}e^{-(1+\delta)j}\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\\ \inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\in[j,j+1]\end{subarray}}\left\lVert\mathbf{v}+\bm{\psi}_{n}(\ell)\right\rVert e^{-\mathbf{c}_{2}\cdot(\mathbf{v}+\bm{\psi}_{n}(\ell))}$
$\ll x^{-\varepsilon/10}\max(1,\min(\ell,n-\ell))^{-C_{2}\left\lVert\mathbf{c}_{2}\right\rVert^{2}}$
$\hskip 85.35826pt\times\sum_{j=0}^{\infty}e^{-(1+\delta)j}\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\\ \inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\in[j,j+1]\end{subarray}}\left\lVert\mathbf{v}\right\rVert e^{-\mathbf{c}_{2}\cdot\mathbf{v}},$
where the last step follows from the definition of $\bm{\psi}_{n}(\ell)$ and $\bm{\psi}_{n}(\ell)\ll\log n\asymp\log x$. By elementary geometry and properties of $\mathbb{K}_{\alpha}(-\mathbf{c}_{2})$, for $j\geqslant 0$,
$\sum_{\begin{subarray}{c}\mathbf{v}\in\mathbb{Z}^{d}\cap\mathbb{K}_{\alpha}(-\mathbf{c}_{2})\\ \inf_{\mathbf{y}\in(\mathbb{K}_{\alpha}(-\mathbf{c}_{2}))^{\mathrm{c}}}\left(\mathbf{c}_{2}\cdot\left(\mathbf{y}-\mathbf{v}\right)\right)\in[j,j+1]\end{subarray}}\left\lVert\mathbf{v}\right\rVert e^{-\mathbf{c}_{2}\cdot\mathbf{v}}\ll e^{-j}.$
Together we obtain
$\mathbb{E}[M_{x,\ell}]$ $\ll x^{-\varepsilon/10}\max(1,\min(\ell,n-\ell))^{-C_{2}\left\lVert\mathbf{c}_{2}\right\rVert^{2}}\sum_{j=0}^{\infty}e^{-\delta j}$
$\ll x^{-\varepsilon/10}\max(1,\min(\ell,n-\ell))^{-C_{2}\left\lVert\mathbf{c}_{2}\right\rVert^{2}}.$
We thus conclude with the following bound on $\mathbb{E}[M_{x}]$:
$\mathbb{E}[M_{x}]=\sum_{\ell=0}^{n}M_{x,\ell}\ll x^{-\varepsilon/4}+x^{-\varepsilon/10}\sum_{\ell=0}^{n-1}\max(1,\min(\ell,n-\ell))^{-C_{2}\left\lVert\mathbf{c}_{2}\right\rVert^{2}}.$
With $C_{2}$ picked large enough, we have $\mathbb{E}[M_{x}]\ll x^{-\varepsilon/10}$.
The rest of the arguments is standard. Recall the weaker bound of (11). If a particle reaches $B_{x}$ at time ntn-t where $t=O(\log x)$, the probability that one of its descendants is located in $B_{x}$ at time $n$ is $\gg(\log x)^{-d}$. On the other hand, $\mathbb{E}[M_{x}]\ll x^{-\varepsilon/10}=o((\log x)^{-d})$. As a consequence, with probability $o(1)$, there is no particle present in $B_{x}$ before time $n=\widehat{A}_{2}(x)$. This finishes the proof of the lower bound of Theorem 3.
2.5 Proofs of remaining corollaries
2.5.1 Proof of Corollary 4
Denote by $\tau_{x}(r)$ the first passage time of the BRW to a sphere of radius $r>0$ centered at $\mathbf{x}=(x,\mathbf{0})$. Note that for a fixed $r>0$, the proofs of Theorems 2 and 3 go through.
Using elementary geometry, we have
$\inf_{x\in[n,n+1]}\tau_{x}(1)\geqslant\max(\tau_{n}(2),\tau_{n+1}(2))$ (49)
and
$\sup_{x\in[n,n+1/2]}\tau_{x}(1)\leqslant\min\Big(\tau_{n}\big(\frac{1}{2}\big),\tau_{n+1/2}\big(\frac{1}{2}\big)\Big).$ (50)
Applying Theorems 2 and 3, we have for any fixed $\varepsilon>0$,
$\sum_{n\in\mathbb{N}}\mathbb{P}\bigg(\Big|\frac{\tau_{n}(2)}{n}-\frac{1}{\widehat{c}_{1}}\Big|>\varepsilon\bigg)+\sum_{n\in\mathbb{N}}\mathbb{P}\bigg(\Big|\frac{\tau_{n/2}(1/2)}{n/2}-\frac{1}{\widehat{c}_{1}}\Big|>\varepsilon\bigg)<\infty.$
By the Borel–Cantelli lemma,
$\frac{\tau_{n}(2)}{n}\to\frac{1}{\widehat{c}_{1}}\quad\text{ and }\quad\frac{\tau_{n/2}(1/2)}{n/2}\to\frac{1}{\widehat{c}_{1}}\quad\text{a.s.}$
Applying (49) and (50) then yields that almost surely,
$\limsup_{x\to\infty}\frac{\tau_{x}(1)}{x}\leqslant\frac{1}{\widehat{c}_{1}}\leqslant\liminf_{x\to\infty}\frac{\tau_{x}(1)}{x}.$
This proves Corollary 4.
2.5.2 Proof of Corollary 5
Denote by $E_{n}$ the event that the range
$R_{n}=\{\bm{\eta}_{v,n}(n):v\in V_{n}\}$
is not $(1/n)$-dense in $(n/C_{2})B_{\mathbf{0}}$ (meaning that there exists $\mathbf{y}\in(n/C_{2})B_{\mathbf{0}}$ with $(\mathbf{y}+(1/n)B_{\mathbf{0}})\cap R_{n}=\emptyset$). To bound $\mathbb{P}(E_{n})$ from above, we need a slight modification of Theorem 3 that applies for balls of radii $1/n$ instead of a fixed radius. By performing the same analysis in Section 2.4.3, we have the upper bound
$\tau_{x,1/n}\leqslant\frac{x}{\widehat{c}_{1}}+C_{1}\log\log x+O_{\mathbb{P}}(1),$
where $\tau_{x,1/n}$ denotes the first passage time to a ball of radius $1/n$ centered at $(x,\mathbf{0})$ and $C_{1}>0$. We need to show that $\widehat{c}_{1}$ is large in all directions. For a $d\times d$ orthogonal matrix $\Theta$, denote by $\widehat{c}_{1}(\Theta)$ the $\widehat{c}_{1}$ corresponding to the BRW with jump distribution $\Theta{\bm{\xi}}$. By assumption (A5) and our assumption that $\phi_{\bm{\xi}}$ is well-defined in a neighborhood of $\mathbf{0}$, the domain of $I$ contains a neighborhood of $\mathbf{0}$, and hence
$\inf_{\Theta}\widehat{c}_{1}(\Theta)>0.$ (51)
Applying (51) gives that there exists $C_{2}>0$, so that $\tau_{x,1/n,\bm{\theta}}\leqslant(C_{2}/2)x+O_{\mathbb{P}}(1)$ in all directions $\bm{\theta}\in\mathbb{S}^{d-1}$. Therefore, by applying Theorem 2,
$\mathbb{P}(E_{n})\leqslant\sum_{j=1}^{n/C_{2}}Cj^{d-1}\sup_{\bm{\theta}\in\mathbb{S}^{d-1}}\mathbb{P}(\tau_{j,1/n,\bm{\theta}}\geqslant n)\leqslant\sum_{j=1}^{n/C_{2}}Cj^{d-1}e^{-n/C_{3}}\ll e^{-n/C_{4}}$
for some $C_{3},C_{4}>0$. Thus $\sum_{n}\mathbb{P}(E_{n})<\infty$. The remaining arguments follow in the same way as Corollary 2.5 of [oz2023].
Remark 3.
We expect that using a similar argument, one can prove uniform versions of (4) and (7) under stronger assumptions on $I$. For instance,
$\sup_{\Theta\in\mathbb{S}^{d-1}}\frac{1}{\log x}\Big|\tau_{\Theta\mathbf{x}}-\Big(\frac{x}{\widehat{c}_{1}(\Theta)}+\frac{d+2}{2\widehat{c}_{1}(\Theta)\partial_{x_{1}}I_{\Theta}(\widehat{c}_{1}(\Theta),\mathbf{0})}\,\log x\Big)\Big|\to 0\quad\text{in probability},$
where $I_{\Theta}$ is the large deviation rate function of $\Theta{\bm{\xi}}$.
3 Extensions and applications
3.1 The delayed branching model
In this section, we revisit the applications of BRW in modeling polymer networks. A polymer network is a three-dimensional disordered structure consisting of long sequences of monomers, called polymer chains, connected via chemical bonds. It turns out that in the study of polymers, random walks provide a decent statistical approximation to polymer chains [flory1960elasticity]. In turn, any polymer chain can be bonded to another chain at a certain site (e.g. a monomer) also by chemical bonds called cross-links.
The resulting network topology is generally random. A chain segment is a connected segment of monomers inside a polymer chain with the property that each of its two ends is connected, via cross-links, to another polymer chain (not necessarily the same one), as mentioned earlier. Chain segments have varying (random) lengths.
As it turns out, macroscopic mechanical properties, such as elasticity and fracture toughness, are largely determined by the distribution of shortest paths (SP) between distant monomers [yin2020topological]. Understanding the statistics of these SP distributions is, therefore, central to predicting network behavior under deformation.
The FPT for the BRW to reach a unit ball at a distance of $x$ can approximate the SP between monomers placed at a distance $x$ from each other in a polymer network [zhang2024modeling]. The intuition can be appreciated from the picture in Figure 4. Locally, the polymer network at a given cross-link is approximated by the branching mechanism of the BRW.
In physical experiments, obtaining SP statistics is infeasible. However, extracting SP statistics from computational simulations of polymer networks using coarse-grained molecular dynamics (CGMD) has been shown promising [mohanty2025strength, yin2020topological, yin2024network]. Now, even though the BRW intuition is local, as argued earlier, and the SP distribution is a global object (involving long distances relative to the size of a single monomer), it was shown based on comparisons to CGMD simulations in [zhang2024modeling], that the SP statistics can be reproduced by FPT statistics of a small variation of the classical BRW model (the reason of which will be explained below), which we refer to as the delayed branching model. The reason for this small variation is to model explicitly the cross-link representing the segment that joins any two polymer chains.
Consider the classical BRW model satisfying $p_{3}>0$ and $p_{0}+p_{1}+p_{3}=1$ (i.e., only branching instances are into three particles) with a spherically symmetric jump distribution (for simplicity). The choice of branching into three particles is to be consistent with the real polymer network being modeled, where branching represents the cross-link (bond) between two polymer chains. The first particle would correspond to the continuation of the first polymer chain beyond the point of cross-link whereas, the other two particles would correspond to the segments of the second polymer chain (see [zhang2024modeling] for a detailed description).
The classical BRW assumes that the three particles emerge from the same point. However, the more realistic picture states that the first and second chains are displaced at the point of branching due to the presence of the cross-link. The delayed branching model is a modification of such a BRW process, where at each branching event (in which the particle was supposed to give rise to three offsprings), a particle branches into two, and one of the two descendants branches again into two particles in the next time step. [^12] A particle may die in each of the two branching sub-events. That is, the branching event (into three particles) is broken into two pieces, one of which is delayed by one step. Thus, each branching event consists of two sub-events, which we call types I and II respectively. Here, type I branching corresponds to the branching of the parent, and type II branching corresponds to the branching of one of the two children born in the type I sub-event.
Figure 4: BRW tree representations of polymer networks in the classical and the delayed branching regimes. The red path from AA to $A^{\prime\prime}$ represents the first polymer chain and the blue path from $B$ to $B^{\prime\prime}$ represents the second chain that cross-links with the first. In the BRW tree genealogy, the particles $A^{\prime\prime}$, $B$, and $B^{\prime\prime}$ are the descendants of the particle starting from $A$.
We first derive the analogue for $\rho$ (the parameter governing the growth of the number of particles) in the classical BRW. By construction, the expected number $\tilde{N}_{n}$ of particles in the delayed branching model at time $n$ satisfies
$\tilde{N}_{n+1}=p_{1}\tilde{N}_{n}+p_{3}\tilde{N}_{n}+2p_{3}\tilde{N}_{n-1}(1-p_{0}).$ (52)
The characteristic polynomial of the recursive equation (52) is given by
$x^{2}-(p_{1}+p_{3})x-2p_{3}(1-p_{0}).$ (53)
Denote the unique positive zero of (53) by
$\tilde{\rho}=\frac{1-p_{0}}{2}+\sqrt{\frac{(1-p_{0})^{2}}{4}+2p_{3}(1-p_{0})}.$ (54)
In particular, $\tilde{N}_{n}\asymp\tilde{\rho}^{n}$.
Let $\widetilde{c}_{1}$ satisfy $I(\widetilde{c}_{1})=\log\tilde{\rho}$ where the rate function $I$ is defined in (1), and $\widetilde{c}_{2}=I^{\prime}(\widetilde{c}_{1})$. We define $\widetilde{A}(x)$ similarly as in (3) with $c_{1},\,c_{2}$ replaced by $\widetilde{c}_{1},\,\widetilde{c}_{2}$. The parameters $p_{3}$ and $p_{0}$ are denoted by $\widetilde{\kappa}$ and $\widetilde{\nu}$ in [zhang2024modeling] respectively, where $\widetilde{\kappa}$ represents the branching rate of the BRW and $\widetilde{\nu}$ represents the death rate of the branches to account for the finite length of the polymer chains in the CGMD simulation. Let $\widetilde{\tau}_{x}$ be the first passage time to $B_{x}$.
Theorem 18.
Assume (A3) and (A4) with $\rho$ replaced by $\tilde{\rho}$. Conditioned upon survival, we have the asymptotic
$\widetilde{\tau}_{x}=\widetilde{A}(x)+O_{\mathbb{P}}(\log\log x)=\frac{x}{\widetilde{c}_{1}}+\frac{d+2}{2\widetilde{c}_{2}\widetilde{c}_{1}}\,\log x+O_{\mathbb{P}}(\log\log x).$ (55)
In other words, the collection $\{(\widetilde{\tau}_{x}-\widetilde{A}(x))/\log\log x\}_{x>0}$ is tight.
Theorem 18 partially resolves Conjecture 1 of [zhang2024modeling]. We sketch its proof below, which will be very similar to the proof of Theorem 1. A key ingredient in the proof of Theorem 1 is the analysis of the expected number $N_{n}$ of particles present at time $n$, conditioned upon survival. If the branching events are independent across layers, then $N_{n}=\rho^{n}$. On the other hand, for the delayed branching BRW, due to the two-step branching regime, the branching events are not independent. The following two lemmas perform a similar analysis related to the number of particles for the delayed branching model. In particular, the analogue for our expected number of descendants for each particle, $\rho$, becomes $\tilde{\rho}$ for the delayed branching model.
Lemma 19 (Lemma 5 of [zhang2024modeling]).
Conditioned upon survival, the expected number of particles $N_{n}$ at time $n$ satisfies $N_{n}\asymp\tilde{\rho}^{n}$.
Lemma 20.
For $1\leqslant s\leqslant n$, let
$N_{n,s}:=\mathbb{E}[\#\{(v,w)\in V_{n}^{2}:v\sim_{s}w\}]$
be the expected number of pairs of particles at level $n$ whose distance in the genealogical tree is equal to $2s$. Conditioned upon survival, $N_{n,s}\asymp\tilde{\rho}^{n+s}$ uniformly in $1\leqslant s\leqslant n$.
Proof.
By Lemma 19, the expected number of choices of the common ancestor $u$ of $v,w$ is $N_{n-s}\asymp\tilde{\rho}^{n-s}$. We now fix $u$. Let $S_{u}$ denote the survival event of the underlying branching process initiated from $u$. Let $N_{u,s}$ and $N^{\prime}_{u,s}$ denote the number of particles initiated from $u$ in $s$ steps conditioned on $S$ and $S_{u}$ respectively. By a coupling argument, if $u$ initiates a branching sub-event of type I,
$\mathbb{E}[N_{u,s}{\mathbbm{1}}_{S}]\geqslant\mathbb{E}[N^{\prime}_{u,s}{\mathbbm{1}}_{S_{u}}]\asymp N_{s}\asymp\tilde{\rho}^{s}.$
Moreover, $\mathbb{E}[N_{u,s}{\mathbbm{1}}_{S}]$ can be controlled by the expected number of particles without conditioning upon survival, which is $\asymp\tilde{\rho}^{s}$. Otherwise, if $u$ initiates a branching sub-event of type II, an extra factor of $2$ exists but will be absorbed into the asymptotics. Therefore we conclude that $\mathbb{E}[N_{u,s}\mid S]\asymp\tilde{\rho}^{s}$ (that $\mathbb{P}(S)\in(0,1)$ has been proven in Lemma 5 of [zhang2024modeling]), and hence $N_{n,s}\asymp\tilde{\rho}^{n-s}(\tilde{\rho}^{s})^{2}=\tilde{\rho}^{n+s}$. ∎
We briefly identify the modifications required for the proof of Theorem 18. Note that the only difference between the classical BRW and delayed branching BRW is the branching structure, i.e., the underlying tree that describes the genealogy of the particles. Given that Theorem 2 holds, the proof of Theorem 1 depends on such structure only through (17) and (18). The corresponding modifications can be made by replacing $\rho$ by $\tilde{\rho}$ in view of Lemmas 19 and 20. The proof of Theorem 2 depends on large deviation estimates for the maximum of delayed branching BRW in (28). Nevertheless, this applies by performing the same proof as in [gantert2018large] by comparing against the maximum of independent random walks (Theorems 3.1 and 5.2 therein). The other changes needed in Theorem 2 are minor and will be omitted for brevity, as the constants can always be adjusted to take into account the delayed branching property. Otherwise, the proof of Theorem 18 goes through almost verbatim.
The same arguments apply for non-spherically symmetric jumps, where we instead assume (A1), (A2), (A5), and (A6) and replace Theorem 1 by Theorem 3.
3.2 Finer asymptotics for non-spherically symmetric jumps
We discuss a few cases with non-spherically symmetric jumps where the asymptotic (7) can be improved. The only instance where we explicitly required spherical symmetry of the law of $\bm{\xi}$} for the proof of Theorem 1 lies in the proof of Proposition 10, where we apply the change of measure (66). More precisely, applying the exponential tilt only in the $x$-coordinate in (66) does not guarantee that the random variable $\bm{\xi}$} is centered in other coordinates under the new measure; cf. (72).
In certain nice cases, the non-spherically symmetric case can be analyzed using the same approach. Let $I$ be the large deviation rate function for the jump $\bm{\xi}$} and $\widehat{c}_{1}$ be such that $I(\widehat{c}_{1},\mathbf{0})=\log\rho$. One may also define the constant $c_{1}$ that arose in assumption (A4) for a non-spherically symmetric jump $\bm{\xi}$}, through the relation $I^{(1)}(c_{1})=\log\rho$ where $I^{(1)}$ is the large deviation rate function for the first marginal of $\bm{\xi}$}. If $\widehat{c}_{1}=c_{1}$, we can show using essentially the same proof of Theorem 1 that
$\tau_{x}=\frac{x}{\widehat{c}_{1}}+\frac{d+2}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\,\log x+O_{\mathbb{P}}(\log\log x).$
The following proposition gives a necessary and sufficient condition for this to happen for all $\rho>1$.
Proposition 21.
It holds that $\widehat{c}_{1}=c_{1}$ for all $\rho>1$ if and only if $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$ for all $2\leqslant i\leqslant d$, where $\xi^{(i)}$} denotes the $i$-th coordinate of the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$}.
Proof.
Since $I^{(1)}$ is strictly increasing on its domain intersecting $(0,\infty)$, an equivalent condition for $\widehat{c}_{1}=c_{1}$ for all $\rho$ is that $I^{(1)}(c_{1})=I(c_{1},\mathbf{0})$ for all $c_{1}$. By definition, this amounts to
$\sup_{\lambda_{1}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\lambda_{1}{\xi^{(1)}}}]\right)=\sup_{\lambda_{1},\dots,\lambda_{d}\in\mathbb{R}}\left(\lambda_{1}c_{1}-\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]\right).$
Since the range of $I^{\prime}$ is $\mathbb{R}_{+}$, this is equivalent to
$\forall\lambda_{1}\in\mathbb{R},\penalty 10000\ \inf_{\lambda_{2},\dots,\lambda_{d}\in\mathbb{R}}\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]=\mathbb{E}[e^{\lambda_{1}{\xi^{(1)}}}].$
By taking partial derivatives with respect to $\lambda_{i}$ at $\lambda_{i}=0,\penalty 10000\ 2\leqslant i\leqslant d$, we obtain that for each $2\leqslant i\leqslant d,\penalty 10000\ \mathbb{E}[{\xi^{(i)}}e^{\lambda_{1}{\xi^{(1)}}}]=0$. Since this holds for all $\lambda_{1}\in\mathbb{R}$, by the uniqueness of the Laplace transform, we must have $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$. The reverse direction is obvious by Jensen’s inequality. ∎
There is yet a different method to deal with distributions on $\mathbb{R}^{d}$ that can be transformed into a spherically symmetric distribution under an invertible linear transformation of space. Suppose that ${\bm{\xi}}\overset{\mathrm{law}}{=}T({\bm{\zeta}})$ for some spherically symmetric distribution $\bm{\zeta}$} and invertible linear map $T$. The FPT for the BRW with jump $\bm{\xi}$} to $B_{x}$ is then equivalent to the FPT for the BRW with jump $\bm{\zeta}$} to $T^{-1}(B_{x})$. Since Theorem 1 holds with the radius of the ball replaced by any constant (with the same proof) and $T$ is invertible, we may without loss of generality replace $T^{-1}(B_{x})$ by $B_{T^{-1}\mathbf{x}}$, where $\mathbf{x}=(x,\mathbf{0})$. This leads to
$\tau_{x}=\frac{\left\lVert T^{-1}\mathbf{x}\right\rVert}{\widehat{c}_{1}({\bm{\zeta}})}+\frac{d+2}{2\widehat{c}_{1}({\bm{\zeta}})\partial_{x_{1}}I_{\bm{\zeta}}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0})}\,\log x+O_{\mathbb{P}}(\log\log x).$ (56)
Summarizing the discussions above, we arrive at the following.
Theorem 22.
Assume assumption (A1). Suppose that either of the following holds:
  1. (i)
    the law of ${\bm{\xi}}=({\xi^{(1)}},\dots,{\xi^{(d)}})$ satisfies $\mathbb{E}[{\xi^{(i)}}\mid{\xi^{(1)}}]=0$ a.s. for all $2\leqslant j\leqslant d$, and satisfies (A2) and (A4);
  2. (ii)
    there exists an invertible linear transformation $T$ on $\mathbb{R}^{d}$ with ${\bm{\xi}}\overset{\mathrm{law}}{=}T({\bm{\zeta}})$ for some spherically symmetric distribution $\bm{\zeta}$} satisfying (A2)–(A4).
Then the first passage time $\tau_{x}$ has the asymptotic
$\tau_{x}=\frac{x}{\widehat{c}_{1}}+\frac{d+2}{2\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\,\log x+O_{\mathbb{P}}(\log\log x).$ (57)
Proof.
By comparing (56) and (57), it suffices to show for case (ii) that
$\frac{x}{\widehat{c}_{1}({\bm{\xi}})}=\frac{\left\lVert T^{-1}\mathbf{x}\right\rVert}{\widehat{c}_{1}({\bm{\zeta}})}$ (58)
and that
$\widehat{c}_{1}({\bm{\xi}})\partial_{x_{1}}I_{\bm{\xi}}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})=\widehat{c}_{1}({\bm{\zeta}})\partial_{x_{1}}I_{\bm{\zeta}}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0}).$ (59)
Let $\mathbf{e}_{1}=(1,\mathbf{0})\in\mathbb{R}^{d}$. By definition (2) and since $T$ is invertible, $I_{\bm{\xi}}(\mathbf{x})=I_{\bm{\zeta}}(T^{-1}\mathbf{x})$. By spherical symmetry of $\bm{\zeta}$},
$I_{\bm{\xi}}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})=I_{\bm{\zeta}}\left(\frac{T^{-1}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0})}{\left\lVert T^{-1}\mathbf{e}_{1}\right\rVert}\right)=I_{\bm{\zeta}}\left(\frac{\left\lVert T^{-1}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0})\right\rVert}{\left\lVert T^{-1}\mathbf{e}_{1}\right\rVert}\right)=I_{\bm{\zeta}}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0}),$
proving (58). Denote the radial component of $I_{\bm{\zeta}}$ by $f$. By applying the multivariate chain rule, we have
$\partial_{x_{1}}I_{\bm{\xi}}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})$ $=\nabla I_{\bm{\zeta}}(T^{-1}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0}))\cdot\partial_{x_{1}}T^{-1}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})$
$=f^{\prime}\left(\left\lVert T^{-1}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})\right\rVert\right)\frac{T^{-1}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})}{\left\lVert T^{-1}(\widehat{c}_{1}({\bm{\xi}}),\mathbf{0})\right\rVert}\cdot T^{-1}\mathbf{e}_{1}$
$=f^{\prime}\left(\widehat{c}_{1}({\bm{\zeta}})\right)\frac{\widehat{c}_{1}({\bm{\xi}})\left\lVert T^{-1}\mathbf{e}_{1}\right\rVert^{2}}{\widehat{c}_{1}({\bm{\zeta}})}$
$=\partial_{x_{1}}I_{\bm{\zeta}}(\widehat{c}_{1}({\bm{\zeta}}),\mathbf{0})\left\lVert T^{-1}\mathbf{e}_{1}\right\rVert.$
Combined with (58), this establishes (59). ∎
Condition (i) handles the case where the law of $\bm{\xi}$} is given by a product measure. The condition (ii) above is satisfied, e.g., for centered non-degenerate elliptical distributions. In particular, this applies to non-degenerate Gaussian distributions, which we explicitly compute in the following example.
Example 23.
Suppose that $\bm{\xi}$} is centered Gaussian with an invertible covariance matrix $\Sigma=AA^{\top}$. In this case, ${\bm{\xi}}=A{\bm{\zeta}}$ where $\bm{\zeta}$} is standard Gaussian. In this case, $I_{\bm{\xi}}(\mathbf{x})=\mathbf{x}^{\top}\Sigma^{-1}\mathbf{x}/2$, so that $\widehat{c}_{1}=\sqrt{2\log\rho/(\Sigma^{-1})_{11}}$ and $\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})=\sqrt{2(\log\rho)(\Sigma^{-1})_{11}}$. By Theorem 22, we have
$\tau_{x}=\frac{x}{\sqrt{2\log\rho/(\Sigma^{-1})_{11}}}+\frac{d+2}{4\log\rho}\,\log x+O_{\mathbb{P}}(\log\log x).$
One can check indeed that this also follows from (56), using $\left\lVert A^{-1}\mathbf{x}\right\rVert=x\sqrt{(\Sigma^{-1})_{11}}$ where $\mathbf{x}=(x,\mathbf{0})$.
4 Numerical analysis
In this section, we test Theorem 3 numerically. Since tracking the locations of all particles in a BRW is time-consuming and memory-intensive, we introduce a path-purging algorithm that removes particles that are highly unlikely to have descendants that realize the first passage time. [^13] An efficient algorithm that simulates the tip of one-dimensional BBM has been proposed in [brunet2020generate]. Unfortunately, the arguments therein do not seem to carry over to general BRWs in higher dimensions. A preliminary version of the algorithm for spherically symmetric jumps can be found in Section 2.3 of [zhang2024modeling].
Recall that $I$ denotes the large deviation rate function of the jump distribution $\bm{\xi}$}. In the path-purging algorithm, we purge along the normal vector
$\mathbf{n}=\nabla I(\widehat{c}_{1},\mathbf{0}).$ (60)
That is, we retain particles that have the $q_{c}/2$ largest inner product with $\mathbf{{n}}$, where $q_{c}$ is a prefixed large number. To intuitively justify this, suppose that a particle is found at $\mathbf{x}^{\prime}=(x^{\prime},\mathbf{0})$ where $x^{\prime}<x$. The expected time for the BRW initiated at $\mathbf{x}^{\prime}$ to arrive at $B_{x}$ is around $(x-x^{\prime})/\widehat{c}_{1}$. In order for another particle located at $\mathbf{x}^{\prime}+\mathbf{v}$ to reach $B_{x}$ within roughly the same amount of time, we need
$e^{-(x-x^{\prime})I(\widehat{c}_{1},\mathbf{0})/\widehat{c}_{1}}=e^{-(x-x^{\prime})I(\widehat{c}_{1}(\mathbf{x}-\mathbf{x}^{\prime}-\mathbf{v})/(x-x^{\prime}))/\widehat{c}_{1}}.$
For small $\mathbf{v}$, this holds if and only if $\mathbf{v}\cdot\nabla I(\widehat{c}_{1},\mathbf{0})=0$, independent of the choice of $x^{\prime}$.
If $\bm{\xi}$} is spherically symmetric, or satisfies condition (i) of Theorem 22, the normal vector $\mathbf{n}$ is then parallel to the direction towards the termination site $\mathbf{x}=(x,\mathbf{0})$.
In our implementation, we take $q_{c}=9000$ and consider BRW where particles only branch into three (i.e., $\{i>1:p_{i}>0\}=\{3\}$) in dimension $d=3$. At times, we also consider the delayed branching model from Section 3.1. We compute the first passage times for offspring distributions with $p_{3}\in\{0.05\ell:1\leqslant\ell\leqslant 19\}$ and $p_{1}=1-p_{3}$ (meaning $p_{0}=0$, unless otherwise stated) while varying $x$ and the distribution of $\bm{\xi}$}. We showcase the following two cases:
  • In Section 4.1, we discuss the case where $\bm{\xi}$} is uniformly distributed on $\mathbb{S}^{2}$.
  • In Section 4.2, we deal with Gaussian distributions with a general covariance matrix.
Finally, we note that the main result of [gantert2011asymptotics] shows that there rarely exists an infinite path in a one-dimensional BRW that always stays above the curve $n\mapsto(c_{1}-\varepsilon)n$ for $\varepsilon>0$ small. While this suggests that our path purging algorithm cannot be backed up by theory for a fixed $q_{c}$ as $x\to\infty$, the simulations work effectively well in the regime where $q_{c}$ is quite large compared to $x$.
4.1 Uniform distribution on the sphere
Consider the case when $\bm{\xi}$} is uniformly distributed in $\mathbb{S}^{2}$, i.e., we are in the spherically symmetric case. The statistic that we compare is the coefficient of the linear term in the estimation of $\tau_{x}$ in (4). For any given $p_{3}\in\{0.05\ell:1\leqslant\ell\leqslant 19\}$, in the numerical BRW simulation, we calculate the FPT distribution at different offset distances, $x\in\{0.25L_{x},\,0.5L_{x},\,0.75L_{x},\,L_{x}\}$, where $L_{x}=65.5$. The mean FPT $\mathbb{E}[\tau_{x}]$ is then fit to a function of the form
$\mathbb{E}[\tau_{x}]=\frac{x}{c_{1}}+B\log x+C.$ (61)
Figure 5: The numerical 3D BRW obtained $c_{1}$ compared against the reference calculations across different $p_{3}$ values for the (a) classical BRW and (b) delayed branching BRW with $p_{0}=0.004$. The mean of the FPT at each $x$ is computed from $2000$ samples.
The numerical implementation validates the theoretical predictions to the classical and delayed branching BRW models, as shown in Figure 5. This also indicates that the numerical approximation of path purging does not affect the linear coefficient of the scaling behavior of the FPT distribution, and can be used as a suitable approximation for carrying out the numerical BRW simulations. The error between the theoretical and numerical implementation can be attributed to the limited range of $x$ over which the fit of (61) is executed. Figure 5 suggests that classical model requires a smaller range of $x$ for the FPT evaluation for a higher fidelity estimation of $c_{1}$ in comparison to the delayed branching model.
Figure 6: (a) The FPT distributions at different values of $x$ for the classical BRW. The (b) mean and (c) standard deviation of the SP as a function of $x$, respectively. The data points correspond to $x=0.1L_{x},\,0.15L_{x},\,0.2L_{x},\,0.25L_{x},\,0.5L_{x},\,0.75L_{x},\,L_{x}$ and $p_{3}=0.0856$ for the classical BRW, where $L_{x}=65.5$. The mean of the FPT at each $x$ is computed from $500$ samples.
Figure 6 shows the FPT distribution at different $x$. Figure 6 shows the linear increase in the mean FPT (negligible logarithmic effect), whereas Figure 6 validates the asymptotic in (55) by showing a negligible change in the standard deviation of the FPT distribution with increasing offset distance, $x$.
4.2 Gaussian distributions
Suppose now that the jump $\bm{\xi}$} is centered Gaussian with a positive definite covariance $\Sigma$. We have $\Lambda(\bm{\lambda})=\bm{\lambda}^{\top}\Sigma\bm{\lambda}/2$ and $I(\mathbf{x})=\mathbf{x}^{\top}\Sigma^{-1}\mathbf{x}/2$. In particular, $\mathbf{c}_{2}=\nabla I(x/n,\mathbf{0})=((\Sigma^{-1})_{11}x/n,\dots,(\Sigma^{-1})_{1d}x/n)$ and $\Lambda(\mathbf{c}_{2})=(x/n)^{2}(\Sigma^{-1})_{11}/2$. By (60), the normal vector $\mathbf{n}$ is a constant multiple of $\Sigma^{-1}\mathbf{e}_{1}$, where $\mathbf{e}_{1}=(1,\mathbf{0})\in\mathbb{R}^{d}$. If $\Sigma$ is diagonal (independent jumps), the normal vector $\mathbf{n}$ is then parallel to the direction of the termination site $\mathbf{x}$. This is the case in our previous work [zhang2024modeling] as opposed to the general case of non-spherically symmetric and dependent jumps that we will discuss in this section.
We look at three specific examples: (i) symmetric and independent (S.I.) jumps (presented earlier [zhang2024modeling]), (ii) non-symmetric and independent (N-S.I.) jumps, and (iii) non-symmetric and dependent (N-S.D.) jumps. The corresponding covariances for the three cases are given by
$\Sigma_{\rm S.I.}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\,\,\,\,\,\,\Sigma_{\rm N\textrm{-}S.I.}=\begin{bmatrix}1&0&0\\ 0&1.5&0\\ 0&0&0.5\end{bmatrix},\,\,\,\,\,\,\Sigma_{\rm N\textrm{-}S.D.}=\begin{bmatrix}1&0.5&0.25\\ 0.5&1.5&0.5\\ 0.25&0.5&0.5\end{bmatrix}$.
We follow the same procedures as Section 4.1 that estimate the statistic $\widehat{c}_{1}$ in (7) through fitting (61). Note that at any given $p_{3}$, the $\widehat{c}_{1}$ is a function of $(\Sigma^{-1})_{11}$. As a result, we expect the $c_{1}$ of the Gaussian BRW with S.I. and N-S.I. jumps to be identical, as confirmed by our numerical results in Figures 7 and 7. We use $p_{0}=0$ for the comparison between the numerical and theoretical results in this section.
Figure 7: The numerically obtained $\widehat{c}_{1}$ compared against the theoretical $\widehat{c}_{1}$ for Gaussian BRW with (a) symmetric and independent (S.I.) jumps, (b) non-symmetric and independent (N-S.I.) jumps, and (c) non-symmetric and dependent (N-S.D.) jumps. The mean of the FPT at each $x$ is computed from $10000$ samples.
In the case of the N-S.D. jumps, the $(\Sigma^{-1})_{11}$ is larger and we expect the $\widehat{c}_{1}$ to be smaller. Figure 7 shows the agreement of the numerical implementation in being able to capture the lower $c_{1}$ at every $p_{3}$. The agreement of the numerical results could be improved by incorporating a larger number of paths for the computation of the mean FPT or by increasing the range of $x$ over which the mean FPT is computed to execute the fit in (61). To ascertain that 500 samples are sufficient to capture the mean, we present the mean FPT $\mathbb{E}[\tau_{x}]$ at different values of $p_{3}$ at $x=65.5$ for the Gaussian BRW with N-S.I. and N-S.D. jumps.
Figure 8: The numerically obtained expected FPT compared against the theoretical asymptotic (7) for Gaussian BRW with (a) non-symmetric and independent (N-S.I.) jumps and (b) non-symmetric and dependent (N-S.D.) jumps. The mean of the FPT is calculated at $x=L_{x}$, where $L_{x}=65.5$, from $10000$ samples.
Figure 8 confirms that the statistics from 500 independent samples are sufficient to capture the mean of the distribution. This confirms that the approximation of $c_{1}$ can be improved by fitting the FPT values over a larger range of $x$.
5 Concluding remarks
In this paper, we study the first passage times of an $\mathbb{R}^{d}$-valued branching random walk (BRW) to a shifted unit ball. The first passage time asymptotics consist of a linear term and a logarithmic correction term, as a function of the Euclidean distance of the target ball from the origin. We discuss extensions of this result to the delayed branching BRW model. As an immediate application, we obtain theoretical predictions of the shortest-path statistics for polymer networks consisting of long chains and random cross-links. We conclude with a conjecture on finer asymptotics of the first passage times.
The first passage times of the branching Brownian motion, the continuous-time sibling of the BRW, possess better tractability due to its connection to the Fisher-KPP equation. It is proven in Section 4.1 of [zhang2024modeling] the finer asymptotic for the first passage times of a standard branching Brownian motion that
$\tau_{x}=\frac{x}{\sqrt{2}}+\frac{d+2}{4}\,\log x+O_{\mathbb{P}}(1),$ (62)
where the $O_{\mathbb{P}}(1)$ is tight. We refer the readers to [zhang2024modeling] for the proof and a detailed discussion. In view also of the tightness result of Theorem 2, it is natural to postulate the following conjecture for BRW. Recall the constant $\widehat{c}_{1}$ from Section 1.2.
Conjecture 1.
Assume (A1)–(A4), or (A1), (A2), (A5), and (A6). Conditioned upon survival, the first passage time for BRW in dimension $d\geqslant 2$ to $B_{x}$ is given by
$\tau_{x}=A(x)+O_{\mathbb{P}}(1)=\frac{x}{\widehat{c}_{1}}+\frac{d+2}{2\,\widehat{c}_{1}\partial_{x_{1}}I(\widehat{c}_{1},\mathbf{0})}\,\log x+O_{\mathbb{P}}(1),$
where the $O_{\mathbb{P}}(1)$ is tight.
Conjecture 1 is also supported by our numerical simulations; see Figure 8.
The main technical obstruction to improving upon $o_{\mathbb{P}}(\log x)$ lies in the application of ballot theorems—the cone $\mathbb{K}$ in Sections 2.4.3 and 2.4.4 cannot be simply replaced by a half-space as certain sums would not converge (if this was the case, Conjecture 1 would hold). One possible approach is to develop bounds for random walks in time-dependent (and possibly non-circular) cones, where $\mathbb{K}=\mathbb{K}_{n}$ may depend on $n$, the number of steps in the random walk. The $o_{\mathbb{P}}(\log x)$ may be improved if one applies such estimates in the case where $\mathbb{K}_{n}\to\mathbb{K}_{\pi/2}(-\mathbf{c}_{2})$ with a certain rate.
Remark 4 (Status of Conjecture 1).
After the first version of this paper was posted, Conjecture 1 was proved in [blanchet2024tightness]. We retain the conjecture here for historical context and refer the readers to Theorems 1 and 2 of [blanchet2024tightness]. Nevertheless, a stronger form of Conjecture 1 — that the $O_{\mathbb{P}}(1)$ term converges in law — remains unsolved.
{acks}
We thank Amir Dembo, Haotian Gu, and Alexandra Stavrianidi for helpful discussions. We also thank two referees for their detailed comments that significantly improved the presentation of this paper.
{funding}
The material in this paper is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, 2229012, 2312204.
Data Availability
The open source code for the branching random walk calculations can be accessed from PolyBranchX.
Large deviation estimates
In this appendix, we collect the deferred proofs of the large deviation estimates (Proposition 10 in Section 2.2.2, Lemma 11 in Section 2.3, and Proposition 17 in Section 2.4.2). We first quote a result on a quantitative version of the central limit theorem for non-lattice distributions.
Lemma 24.
Consider $d\geqslant 2$ and a compact set $K\subseteq\mathbb{R}^{d}$. Suppose that the radial component $R$ of the spherically symmetric random variable $\bm{\xi}$} has exponential moments and $\mathbb{P}(R=0)<1$. Let $\bm{\lambda}\in\mathbb{R}^{d}$ be such that $\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ exists and define a probability measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\bm{\lambda}\cdot\mathbf{x}}/\mathbb{E}_{\mathbb{P}}[e^{\bm{\lambda}\cdot\mathbf{X}}]$ where $\mathbb{P}$ is the law of $\mathbf{X}$. Then the law of $\bm{\xi}$} under $\mathbb{Q}$ satisfies the strong non-lattice property
$\limsup_{|\mathbf{t}|\to\infty}|\mathbb{E}_{\mathbb{Q}}[e^{i\mathbf{t}\cdot{\bm{\xi}}}]|<1.$ (63)
Moreover, consider an i.i.d. sequence $\{\mathbf{X}_{n}\}_{n\geqslant 1}$ with the same distribution as $\bm{\xi}$} under $\mathbb{Q}$ with a positive definite covariance matrix $V$ and suppose that $\mathbb{E}_{\mathbb{Q}}[\left\lVert{\bm{\xi}}\right\rVert^{d+2}]<\infty$. Then
$\sup_{B}\bigg|\mathbb{P}\Big(\frac{1}{\sqrt{n}}\sum_{j=1}^{n}(\mathbf{X}_{j}-\mathbb{E}_{\mathbb{Q}}[\mathbf{X}_{j}])\in B\Big)-h(B)\bigg|=o(n^{-d/2}),$ (64)
where the supremum is taken over all boxes $B\subseteq K$, $h(B)=(1+o(1))\mu_{\mathbf{0},V}(B)$ as $n\to\infty$, and $\mu_{\mathbf{0},V}$ is the Gaussian measure on $\mathbb{R}^{d}$ with mean $\mathbf{0}$ and covariance $V$.
Proof.
Let us first assume that $\mathbb{P}(R=0)=0$. Denote by ${\bm{\Xi}}=(\Xi_{1},\dots,\Xi_{d})$ the uniform distribution on $\mathbb{S}^{d-1}$. Suppose that ${\bm{\xi}}_{1}=R_{1}{\bm{\Xi}}^{(1)}$ and ${\bm{\xi}}_{2}=R_{2}{\bm{\Xi}}^{(2)}$ are i.i.d. copies of $\bm{\xi}$}, and ${\bm{\zeta}}={\bm{\xi}}_{1}+{\bm{\xi}}_{2}$; since each ${\bm{\xi}}_{i}$ is radially symmetric, $R_{i}$ and ${\bm{\Xi}}^{(i)}$ are independent. We first find the conditional density $z_{R_{1},R_{2}}(r,\bm{\theta})$ of $\bm{\zeta}$} under $\mathbb{P}$ given $R_{1}$ and $R_{2}$, where $r$ and $\bm{\theta}$ are the radial and polar components. Observe that the conditional distribution of $\bm{\zeta}$} given $R_{1}$ and $R_{2}$ is spherically symmetric. We have conditionally,
$\mathbb{P}(\left\lVert{\bm{\zeta}}\right\rVert\geqslant r\mid R_{1},R_{2})$ $=\mathbb{P}(\left\lVert(R_{1},\mathbf{0})+R_{2}{\bm{\Xi}}\right\rVert\geqslant r)$
$=\mathbb{P}\left(R_{1}^{2}+2R_{1}R_{2}\Xi_{1}+R_{2}^{2}\geqslant r^{2}\right)=\mathbb{P}\left(\Xi_{1}\geqslant\frac{r^{2}-R_{1}^{2}-R_{2}^{2}}{2R_{1}R_{2}}\right).$
By Theorem 2.10 of [fang2018symmetric], for $y\in[-1,1]$ and some $C(d)\geqslant 0$,
$\mathbb{P}\left(\Xi_{1}\geqslant y\right)=C(d)\int_{-1}^{y}(1-t^{2})^{(d-3)/2}\mathrm{d}t.$
Altogether it follows that for some $C(d)>0$,
$z_{R_{1},R_{2}}(r,\bm{\theta})=C(d)r\left(1-\left(\frac{r^{2}-R_{1}^{2}-R_{2}^{2}}{2R_{1}R_{2}}\right)^{2}\right)^{\frac{d-3}{2}},\penalty 10000\ r\in[|R_{1}-R_{2}|,R_{1}+R_{2}],\penalty 10000\ \bm{\theta}\in\mathbb{S}^{d-1}.$
Therefore, the conditional law of $\bm{\zeta}$} under $\mathbb{P}$, and hence also the unconditional law under $\mathbb{Q}$, has a density in $\mathbb{R}^{d}$. If $\mathbb{P}(R=0)>0$, the laws have densities except for an atom at zero, and the densities are non-trivial since $\mathbb{P}(R=0)<1$. By Lemma 4 in Section XV.4 of [feller1991introduction], (63) holds with $\bm{\zeta}$} in place of $\bm{\xi}$}, and hence (63) holds since $\bm{\zeta}$} is a sum of two i.i.d. copies of $\bm{\xi}$}. The final claim follows from Theorem 2(b) of [bhattacharya1978validity], by noting that the sum in (1.14) therein is $o(1)$ as $n\to\infty$ uniformly on compact sets. ∎
Proof of Proposition 10.
Let $\Lambda(\lambda)=\log\phi_{\xi^{(1)}}(\lambda)=\log\mathbb{E}[e^{\lambda{\xi^{(1)}}}]$. It is clear from (A2) that $\xi^{(1)}$} is non-lattice. The Bahadur–Rao theorem (Theorem 3.7.4 of [dembo2009large]) then implies that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(F_{n}\geqslant m_{n}+c)\asymp\frac{e^{-nI((m_{n}+c)/n)}}{\sqrt{n}}\asymp\frac{e^{-nI(c_{1})-(c-3\log n/2)I^{\prime}(c_{1})}}{\sqrt{n}}=n\rho^{-n}e^{-c_{2}c}.$ (65)
By definition $c_{2}=I^{\prime}(c_{1})$, the constant $c_{2}>0$ is such that the measure $\mathbb{Q}$ defined by
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}(x):=e^{c_{2}x-\Lambda(c_{2})}$ (66)
satisfies that under $\mathbb{Q}$, $\{X_{i}\}_{i\in\mathbb{N}}$ are i.i.d. with mean $c_{1}$. The random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-c_{1}$. Define $\widetilde{F}_{n}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(F_{n}\geqslant m_{n}+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n)))$
$=e^{n\Lambda(c_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-c_{2}F_{n}}{\mathbbm{1}}_{\{F_{n}\geqslant m_{n}+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\}}\right]$
$\asymp e^{n\Lambda(c_{2})-m_{n}c_{2}}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\widetilde{F}_{n}\in\Big[j-1-\frac{3}{2c_{2}}\,\log n,j-\frac{3}{2c_{2}}\,\log n\Big),\,\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\right)$
$=\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big)\times B_{\mathbf{0}}\Big(\frac{u(n)}{\sqrt{n}}\Big)\right).$
Note that for $j\geqslant\sqrt{n}$, we may bound the probability from above by $1$. For $j\leqslant\sqrt{n}$, the sets
$E_{j,n}:=\Big[\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big)\times B_{\mathbf{0}}\Big(\frac{u(n)}{\sqrt{n}}\Big)$
are bounded as $n\to\infty$. Therefore, by Lemma 24, with $\mu_{\mathbf{0},V}$ denoting the centered Gaussian measure with covariance matrix $V$ on $\mathbb{R}^{d}$,
$\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in E_{j,n}\right)=(1+o(1))\mu_{\mathbf{0},V}\left(E_{j,n}\right)\asymp u(n)^{d-1}n^{-d/2},$ (67)
where the last step follows from the multivariate Gaussian density formula and that the positive definite matrix $V$ depends only on the law of $\bm{\xi}$}. Altogether this yields
$\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(u(n))\mid F_{n}\geqslant m_{n})$ $\ll(n\rho^{-n}e^{-c_{2}c})^{-1}\rho^{-n}n^{3/2}\sum_{j=c+1}^{\infty}e^{-c_{2}j}u(n)^{d-1}n^{-d/2}$
$\ll u(n)^{d-1}n^{-(d-1)/2}.$
This establishes (20).
Using a similar approach, we see that
$\mathbb{P}(F_{n}\in[m_{n}+c,m_{n}+c+a(n)])\asymp n\rho^{-n}e^{-c_{2}c}$
and that
$\hskip 14.22636pt\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1),\penalty 10000\ F_{n}\in[m_{n}+c,m_{n}+c+a(n)])$
$\asymp\rho^{-n}n^{3/2}\sum_{j=c+1}^{c+a(n)}e^{-c_{2}j}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}},\frac{j-\frac{3}{2c_{2}}\,\log n}{\sqrt{n}}\Big)\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$
$\gg\rho^{-n}n^{-(d-3)/2}e^{-c_{2}c}.$
This proves (21). [^14] We have implicitly used that $a(n)\geqslant 1$, but the general case $a(n)\gg 1$ proceeds with the same proof by decomposing $m_{n}+c,m_{n}+c+a(n)]$ into sub-intervals of smaller length than one. The same computation also yields
$\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{0}}(K\sqrt{n}),\penalty 10000\ F_{n}\geqslant m_{n}+c)$ $\ll\rho^{-n}n^{3/2}n^{-1/2}=n\rho^{-n},$
where the asymptotic constant does not depend on $K$. Together with (65) completes the proof of (22). ∎
Proof of Lemma 11.
We use a similar change of measure argument as in the proof of Proposition 10. Denote by $I$ the large deviation rate function for the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$} (defined in (2)). For $\widehat{\mathbf{c}}_{2}:=\nabla I({\mathbf{z}}/(Ck))$, we define the tilted measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\widehat{\mathbf{c}}_{2}\cdot\mathbf{x}-\Lambda(\widehat{\mathbf{c}}_{2})}$, where we recall $\Lambda=\log\phi_{\bm{\xi}}$ is the log-moment generating function. Under $\mathbb{Q}$, each ${\bm{\xi}}_{i}$ is i.i.d. with mean ${\mathbf{z}}/(Ck)$. It follows that
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})$ $=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})}\mathbb{E}_{\mathbb{Q}}[e^{-\widehat{\mathbf{c}}_{2}\cdot(F_{Ck},\mathbf{J}_{Ck})}{\mathbbm{1}}_{\{(F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}}\}}]$
$\gg e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{F_{Ck}}{\sqrt{Ck}},\frac{\mathbf{J}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{{\mathbf{z}}}}{\sqrt{Ck}}\right)$
$=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{\widetilde{F}_{Ck}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{J}}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right),$ (68)
where each $(\widetilde{X_{i}},\widetilde{\mathbf{Y}}_{i})=(X_{i},\mathbf{Y}_{i})-{\mathbf{z}}/(Ck)$ is a centered random variable and $(\widetilde{F}_{Ck},\widetilde{\mathbf{J}}_{Ck})=\sum_{j=1}^{Ck}(\widetilde{X_{i}},\widetilde{\mathbf{Y}}_{i})$. Using definition of $\widehat{\mathbf{c}}_{2}$, we have
$\Lambda(\widehat{\mathbf{c}}_{2})=\nabla I\Big(\frac{{\mathbf{z}}}{Ck}\Big)\cdot\frac{{\mathbf{z}}}{Ck}-I\Big(\frac{{\mathbf{z}}}{Ck}\Big).$ (69)
In addition, Lemma 24 yields [^15] Lemma 24 is stated for radially symmetric distributions, but the part that derives (64) from (63) does not depend on the symmetry and applies to general non-lattice distributions. Note that the non-lattice condition (A5) is preserved under exponential tilts.
$\mathbb{Q}\left(\left(\frac{\widetilde{F}_{Ck}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{J}}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right)\gg(Ck)^{-d/2}.$ (70)
Combining (68), (69), and (70) yields that
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})\gg e^{-CkI({\mathbf{z}}/(Ck))}(Ck)^{-d/2}.$
Recalling that $I$ is convex, $C^{1}$ on its domain, and $I(\mathbf{0})=\nabla I(\mathbf{0})=0$ (by Jensen’s inequality, $\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ is minimized at $\bm{\lambda}=\mathbf{0}$ when $\bm{\xi}$} is centered), we have uniformly for $\mathbf{z}$} with $\left\lVert{\mathbf{z}}\right\rVert\leqslant C_{1}k$,
$CI\Big(\frac{{\mathbf{z}}}{Ck}\Big)\leqslant C\sup_{\left\lVert\mathbf{x}\right\rVert_{2}=C_{1}}I\Big(\frac{\mathbf{x}}{C}\Big)\leqslant C_{1}\sup_{\left\lVert\mathbf{x}\right\rVert_{2}=C_{1}}\left\lVert\nabla I\Big(\frac{\mathbf{x}}{C}\Big)\right\rVert_{2}\leqslant\frac{c_{3}}{2}$
for $C\geqslant C_{2}$ as we pick $C_{2}$ large enough. This concludes the proof. ∎
Proof of Proposition 17.
The Bahadur–Rao theorem (Theorem 3.7.4 of [dembo2009large]) yields that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(F_{n}\geqslant x+c)\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}.$ (71)
Mimicking (66), define the probability measure $\mathbb{Q}$ by
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}(\mathbf{x}):=e^{\mathbf{c}_{2}\cdot\mathbf{x}-\Lambda(\mathbf{c}_{2})}.$ (72)
It follows that $\mathbb{E}_{\mathbb{Q}}[{\bm{\xi}}]=(x/n,\mathbf{0})$. That is, the random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-x/n$. Define $\widetilde{F}_{n}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(F_{n}\geqslant x+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1))$
$=e^{n\Lambda(\mathbf{c}_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-\mathbf{c}_{2}\cdot(F_{n},\mathbf{J}_{n})}{\mathbbm{1}}_{\{F_{n}\geqslant x+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\}}\right]$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\widetilde{F}_{n}\in[j-1,j),\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\right)$
$=e^{-nI(x/n,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big)\times B_{\mathbf{0}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$
$\asymp e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{0})}n^{-d/2},$
where in the last step we apply the same argument that leads to (67) and use that the first coordinate of $\mathbf{c}_{2}$ is positive by our assumption $x/n>\underline{c}>0$. Altogether this yields (41).
Using a similar approach, we see that
$\mathbb{P}(F_{n}\in[x+c,x+c+a(n)])\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}$
and that
$\hskip 14.22636pt\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1),\penalty 10000\ F_{n}\in[x+c,x+c+a(n)])$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{y})}\sum_{j=c+1}^{c+a(n)}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big)\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$
$\gg e^{-nI(x/n,\mathbf{0})-\mathbf{c}_{2}\cdot(c,\mathbf{y})}n^{-d/2}.$
This proves (42). ∎